### v1

# Generalized Distribution Prediction for Asset Returns

Abstract
We present a novel approach for predicting the distribution of asset returns using a quantile-based method with Long Short-Term Memory (LSTM) networks. Our model is designed in two stages: the first focuses on predicting the quantiles of normalized asset returns using asset-specific features, while the second stage incorporates market data to adjust these predictions for broader economic conditions. This results in a generalized model that can be applied across various asset classes, including commodities, cryptocurrencies, as well as synthetic datasets. The predicted quantiles are then converted into full probability distributions through kernel density estimation, allowing for more precise return distribution predictions and inferencing. The LSTM model significantly outperforms a linear quantile regression baseline by 98% and a dense neural network model by over 50%, showcasing its ability to capture complex patterns in financial return distributions across both synthetic and real-world data. By using exclusively asset-class-neutral features, our model achieves robust, generalizable results.
keywords:
Machine Learning, Distribution Prediction, Asset Return Forecasting, Long Short-Term Memory (LSTM) Networks
1 Introduction
Every day, an almost uncountable number of transactions take place in capital markets, with each transaction generally aimed at generating a positive return on investment. If one could predict future financial asset prices accurately, the potential for gains would be nearly limitless. Even achieving a prediction accuracy slightly above 50% could potentially turn someone into a billionaire [35]. However, predicting returns is an extraordinarily difficult problem. Nobel laureate Eugene Fama famously argued that future returns are inherently random and, therefore, practically impossible to predict [11]. This difficulty is further supported by the fact that many funds consistently fail to outperform the S&P 500, a common benchmark for returns [13, 19].
In finance, logarithmic asset returns are often assumed to follow a normal distribution, a famous example is the Black-Scholes model which is used to find the fair value of option derivatives [3]. However, in practice, this assumption frequently does not hold [26]. Returns often display non-normal characteristics, such as skewness or heavy tails. Consequently, models that rely on normality assumptions can lead to suboptimal or even incorrect results [30]. To address this issue, some approaches generalize assumptions or adopt alternative distributions, such as the Student’s T distribution, which accommodates skewness, kurtosis, and heavy tails [1]. More recently, a new approach has emerged, namely to model the distribution directly [2]. This approach offers several key advantages. A more accurate model of the return distribution could improve derivative pricing, inform risk management, and enhance portfolio optimization.
In this paper, we propose a model that uses a distributional approach to predict asset returns. Our work extends the model presented in [2], which predicts the distribution of stock returns. Instead of predicting the return itself, our model predicts the quantiles of the return distribution. Although quantile regression is not a novel concept, we convert these quantiles into a full distribution using kernel density estimation. This enables a more precise estimation of the return distribution over a given time period. The model is trained on a dataset of asset returns and general market variables, ensuring that each feature is asset-class neutral, making the approach applicable across different asset classes.
This paper contributes to the literature by creating a more generalizable distribution prediction model, utilizing only asset-neutral features and allowing for the consideration of multiple return types. The results are compared to a baseline linear quantile regression model and the original approach in [2]. We find that our generalized approach outperforms the baseline by 121% on real-world data, while slightly underperforming by 1% on synthetic data. Compared to the original approach, our model outperforms on both real-world and synthetic data by 50% and 64%, respectively.
2 Literature Review
2.1 Asset Return Prediction
The field of asset return prediction using machine learning and artificial intelligence has expanded rapidly in recent years. The sheer volume of published studies makes it increasingly difficult to conduct a thorough review of the existing literature. As noted by Dessain, the number of academic studies has grown at an exponential rate, creating challenges for researchers attempting to synthesize existing knowledge [8]. He observes that conducting a comprehensive review of the literature on machine learning in finance is "becoming more and more challenging, if feasible at all," due to the rapid influx of studies [8]. Similarly, Strader et al. [31] highlighted the difficulty in comparing studies because of the vast diversity of approaches and benchmarks used across research.
In addition to the sheer volume of publications, the complexity and diversity of machine learning models used in asset return prediction further complicates synthesis. The review by Shah et al. underscores the variety of methods employed, including ARIMA [20, 9, 25], LSTM [28, 34, 21], CNN [5, 15, 17], and hybrid models [32, 24, 14]. This wide range of methodologies makes it challenging to compare results across studies [29]. This issue aligns with Dessain’s observation that conducting a comprehensive review is exceptionally difficult [8].
Another factor that hinders a thorough and consistent literature review is the variation in data used across studies [8]. Differences in asset classes, individual assets, time periods, and time intervals, often chosen arbitrarily, complicate meaningful comparisons between methodologies. To mitigate this, our paper uses publicly available data, ensuring that our results can be easily compared to other studies.
While several studies have attempted to address these challenges [27, 18, 4], they too struggle to keep pace with the rapid rate of publication and face the same issues described above. These efforts are important in their own right, but the complexities of the field continue to pose significant challenges for comprehensive reviews.
2.2 Distribution Prediction
Distribution prediction is a relatively novel approach in the machine learning literature. One of the few papers to focus on this is [2], where the authors propose a model that predicts quantiles of the return distribution rather than the return itself. This approach aligns with the quantile regression framework discussed by Liu [23]. The concept has been further advanced by models like the Conditional Quantile Variational Autoencoder (CQVAE), introduced by Yang et al. [33], which employs neural networks to model conditional quantiles. Together, these studies demonstrate the potential of distribution prediction and highlight its advantages over more traditional methods, which are often constrained by certain assumptions about return distributions.
Despite these advances, there remains significant room for improvement, and this is where our work contributes. While [2] predicts the 22-day ahead quantiles of returns for a specific asset, our model leverages Long Short-Term Memory (LSTM) networks to predict quantiles for an $n$-day period, where $n$ corresponds to both the input sequence length and the forecast window.
Our approach also differs in the type of input features used. The model in [2] incorporates 194 stock-specific characteristics along with macroeconomic variables to predict return quantiles. In contrast, our model employs asset-neutral features, drawing on the generalization techniques from the CQVAE model to ensure robustness and applicability across various asset classes.
3 Methodology
In this study, we employ a two-stage model to predict quantiles of asset returns. The first stage uses only asset-specific data to predict quantiles for normalized returns. The second stage incorporates market data to generate a scaling factor, which is then multiplied by the output from the first stage. This scaling factor adjusts the normalized returns based on broader market conditions. To train the model, we utilize a modified quantile loss function that accommodates multiple observations, in line with the generalized approach. This allows for a more robust evaluation of the model’s predictive capability across different datasets. We validate the methodology using both real-world and synthetic data.
The code used to implement all approaches in this section as well as to evaluate the models is publicly available at https://github.com/izzak98/dist-pred
3.1 Problem Definition
An asset is represented as a matrix of input features $X\in\mathbb{R}^{m\times t}$, where $m$ is the number of features, and $t$ is the number of historical time steps. The matrix $X$ is structured as follows
$X=\begin{bmatrix}x_{1,1}&x_{1,2}&\cdots&x_{1,t}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,t}\\ \vdots&\vdots&\ddots&\vdots\\ x_{m,1}&x_{m,2}&\cdots&x_{m,t}\\ \end{bmatrix}$
Accompanying the asset matrix is the market data for the same $t$ time steps, represented as a matrix $Z\in\mathbb{R}^{g\times t}$, where $g$ is the number of market features, that is
$Z=\begin{bmatrix}z_{1,1}&z_{1,2}&\cdots&z_{1,t}\\ z_{2,1}&z_{2,2}&\cdots&z_{2,t}\\ \vdots&\vdots&\ddots&\vdots\\ z_{g,1}&z_{g,2}&\cdots&z_{g,t}\\ \end{bmatrix}$
Our goal is to predict the distribution of log-returns of the asset for the future time steps from $t+1$ to $t+n$, where $n$ represents both the number of time steps used for the predictions and the forecast horizon. The vector of log returns $\vec{r}\in\mathbb{R}^{n}$ is defined as:
$\vec{r}=\begin{bmatrix}r_{t+1}\\ r_{t+2}\\ \vdots\\ r_{t+n}\\ \end{bmatrix}$
where each $r_{t+k}$ represents the log return at time step $t+k$ for $k=1,\dots,n$.
Given the input features $X$ and $Z$, the objective is to model and predict the distribution of future log returns
$\vec{r}\sim f(X,Z)$
where $f(X,Z)$ represents the predictive model for the log returns based on the asset and market features.
3.2 Model Architecture
To model the distribution of future log returns, we aim to predict the quantiles $\vec{\tau}$ of the return distribution over the prediction horizon. Utilizing kernel density estimation these can later be converted into distributions. The quantile vector is defined as
$\vec{\tau}=\begin{bmatrix}\tau_{1}\\ \tau_{2}\\ \vdots\\ \tau_{\mathcal{T}}\\ \end{bmatrix}$
where $\mathcal{T}$ denotes the number of quantiles, the quantiles used can be found in F.
For this task, we utilize Long Short-Term Memory (LSTM) networks, a type of recurrent neural network (RNN) [16] that are well-suited for sequential data modeling due to their ability to capture long-term dependencies and are less prone to issues such as vanishing gradients. Unlike many other time-series models, LSTMs can flexibly handle variable-length sequences, making them appropriate for our task as we want a variable amount of sequence lengths.
Following the methodology outlined in [2], we adopt a two-stage approach to predict the distribution of future log returns.
3.2.1 Stage 1: Asset-Specific Quantile Prediction
In the first stage, the model focuses on predicting the quantiles of future normalized returns $\vec{\tilde{r}}$ by leveraging the asset-specific features $X$ and the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group. This is done by predicting the quantiles $Q_{\vec{\tilde{r}}}^{\tau}$ of the volatility-normalized returns $\vec{\tilde{r}}$, where
$\vec{\tilde{r}}=\frac{\vec{r}}{\vec{\bar{\sigma}}}$
with
$\vec{r}=\begin{bmatrix}r_{t+1}\\ \vdots\\ r_{t+n}\end{bmatrix}\quad\textrm{and}\quad\vec{\bar{\sigma}}=\begin{bmatrix}\bar{\sigma}_{t+1}\\ \vdots\\ \bar{\sigma}_{t+n}\end{bmatrix}$
Here, $\vec{\bar{\sigma}}$ represents the average volatility of assets in the same group at time step $t+k$, for $k=1,\dots,n$, where $n$ is the prediction horizon.
The average volatility $\bar{\sigma}_{t}$ is calculated as the exponentially weighted moving average (EWMA) of the log returns across all $K$ assets in the group, with a decay factor of 0.94, in line with standard practices and the methodology from [2], formally:
$\sigma_{t}^{2}=\lambda\sigma_{t-1}^{2}+(1-\lambda)r_{t-1}^{2}$
with $\lambda$ being the decay factor, in our case 0.94, and $\bar{\sigma}_{t}$ is calculated as:
$\bar{\sigma}_{t}=\frac{1}{K}\sum_{i=1}^{K}\sigma_{i,t}$
where $\sigma_{i,t}$ is the volatility of asset $i$ at time $t$. We can formally represent this stage as:
$f_{1}(X)=Q_{\vec{\tilde{r}}}^{\tau}$
3.2.2 Stage 2: Market Data Scaling
In the second stage, the model incorporates the broader market data features $Z$ to refine the predictions from the first stage. The market data is used to create a scaling factor $\vec{\tilde{\sigma}}$, which adjusts the asset-specific quantile predictions based on market conditions.
The scaling factor $\vec{\tilde{\sigma}}$ is generated by applying a transformation to the market feature matrix $Z$, capturing the influence of market conditions on future returns. Additionally, the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group is used to further scale the predicted quantiles. This results in the final quantile prediction for the asset’s log returns:
$Q_{\vec{r}}^{\tau}=\vec{\hat{\sigma}}\cdot\vec{\bar{\sigma}}\cdot Q_{\vec{\tilde{r}}}^{\tau}$
[…]
$f_{2}(Z)=\vec{\hat{\sigma}}$
Giving the final model the representation:
$\vec{r}\sim f(X,Z)=f_{\text{KDE}}(f_{1}(X)\cdot\vec{\bar{\sigma}}\cdot f_{2}(Z)$
where $f_{\text{KDE}}$ represents the kernel density estimations, for more details see Appendix D.
3.2.3 Loss Function
Following the methodology of [2], we use a custom loss function based on quantile regression to train the model. The original loss function is designed to minimize the error between the true returns and the predicted quantiles. It is defined as
$L_{\mathcal{T}}=\frac{1}{BK}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i}-\hat{Q}_{r_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i}-\hat{Q}_{\tilde{r}_{i}}(\tau)\right)\right]$ (1)
where $\mathcal{T}$ is the set of quantiles used for the prediction, $B$ is the batch size (i.e., the number of assets in each batch), $K$ is the number of quantiles in the set $\mathcal{T}$, $r_{i}$ is the raw return of the asset, $\tilde{r}_{i}$ is the standardized (volatility-normalized) return of the asset and $\hat{Q}_{r_{i}}(\tau)$ and $\hat{Q}_{\tilde{r}_{i}}(\tau)$ are the predicted quantiles of the raw and standardized returns for quantile $\tau$, respectively Furhtermore, $\rho_{\tau}(\xi)$ is the quantile loss function
$\rho_{\tau}(\xi)=\begin{cases}\tau\cdot\xi,&\text{if }\xi\geq 0\\ (\tau-1)\cdot\xi,&\text{if }\xi<0\end{cases}$
where $\xi$ represents the error (residual) between the true return value and the predicted quantile.
3.2.4 Modified Loss Function
Our proposed model generalizes the model in [2] by generating multiple predictions over a look-ahead window, and we therefore modify the loss function in 1 to account for each future time step $t\in\{1,2,\dots,T\}$ within the prediction horizon. The modified loss function is defined as
$L_{\mathcal{T}}=\frac{1}{BTK}\sum_{\tau\in\mathcal{T}}\sum_{t=1}^{T}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i,t}-\hat{Q}_{\vec{r}_{i,t}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i,t}-\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)\right)\right]$
where $\mathcal{T}$ is the number of future time steps (i.e., the prediction horizon) in the look-ahead window, $r_{i,t}$ and $\tilde{r}_{i,t}$ are the raw and standardized returns of the asset at future time step $t$, and $\hat{Q}_{\vec{r}_{i,t}}(\tau)$ and $\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)$ are the predicted quantiles of the raw and standardized returns for asset $i$ at future time step $t$.
This modification allows us to compute the loss over the entire prediction window, training the model to minimize the quantile loss for each future time step within the horizon. The inclusion of both raw and standardized returns helps ensure that the model learns both absolute and volatility-adjusted relationships.
3.3 Market Data
We utilize a diverse set of financial assets and market data for our analysis. The assets span various categories and markets, ensuring comprehensive coverage. The data spans from February 3, 2000 up to January 1, 2024, sourced from the Yahoo Finance API. All data are normalized using rolling Z-scores, with the window length set as a hyperparameter to be optimized during model training.
3.3.1 Asset Selection
Several asset classes are chosen to create a diverse representation to emphasize the generalizability of the model. An overview can be found in table 1. Four asset classes are chosen: stocks, currencies, commodities, and cryptocurrencies to capture a diverse set of asset classes to test the robustness of the model. Within stocks there are three sub-asset classes for the index the stocks are listed on, S&P 500, Euro Stoxx 50, and Nikkei 225. Full details on the selected assets as well as summary statistics, can be found in Appendix A.
Table 1: Summary of Selected Assets
Asset Class Selection
Stocks (S&P 500) 10 stocks from the S&P 500 index.
Stocks (Euro Stoxx 50) 10 stocks from the Euro Stoxx 50 index.
Stocks (Nikkei 225) 10 stocks from the Nikkei 225 index.
Currencies 10 currency pairs.
Commodities 10 commodities.
Cryptocurrencies 10 cryptocurrencies.
3.3.2 Feature Engineering
Additional features are computed for each asset to enhance the model’s ability to capture various market dynamics. Table 2 shows an overview of the engineered features. These features span returns, volatility metrics, moving averages, technical indicators, risk metrics, and the asset class to which the asset belongs. These features are common practice in technical analysis and are widely used across studies [7, 10, 6, 12].
Table 2: Overview of Engineered Features
Feature Category Features
Returns 2-day, 5-day, 22-day returns, cumulative returns, log returns.
Volatility Metrics 2-day, 5-day, 22-day volatility, skewness, kurtosis.
Moving Averages 2-day, 5-day, 22-day simple and exponential moving averages (SMA, EMA).
Technical Indicators RSI, MACD, Bollinger Bands, Stochastic Oscillator, VWAP.
Risk Metrics 2-day, 5-day, 22-day Sharpe Ratio.
Categorical Features Asset category.
3.3.3 Market Data
In addition to the assets, general market data is incorporated to capture broader trends and enhance the model’s ability to account for macroeconomic factors. An overview of the market data sources can be found in Table 8. This data includes stock indices, commodities, other indices, and treasury yields. For more details on market data, see Appendix B.
Table 3: Overview of Market Data Sources
Data Category Data Source
Stock Indices S&P 500, NASDAQ, Dow Jones, Russell 2000, Euro Stoxx 50, Nikkei 225.
Commodities Crude oil, gold, silver futures.
Other Indices Baltic Dry Index, VIX, US Dollar Index, Corporate Bond Index (Bloomberg Barclays).
Treasury Yields 10-year and 3-month Treasury yields.
3.3.4 Data Split
The data is split into training, validation, and test sets based on the time periods as follows: Training set: All available data up to the end of 2017. Validation set: Data from 2018 to the end of 2019. Test set: Data from 2019 to 2024, which includes the COVID-19 pandemic and the 2022 bear market. To create varying sequence lengths, each asset’s data is randomly divided into sequences with lengths sampled uniformly from the range $[1,30]$ days.
3.4 Synthetic Data
[…]
$r_{\text{synthetic}}=\mathcal{D}\cdot\ddot{\sigma}+\ddot{\mu}$
We generate the synthetic return data using the following distributions:
Normal Distribution
$\mathcal{D}\sim\mathcal{N}(0,1)$
where $\mathcal{N}(0,1)$ represents a standard normal distribution with mean 0 and standard deviation 1.
Gamma Distribution
$\mathcal{D}\sim\Gamma(2,1)$
where $\Gamma(2,1)$ represents a gamma distribution with shape parameter 2 and scale parameter 1. The shape and scale parameters were selected to provide variability while covering a realistic range of potential returns. The distribution is then standardized to have a mean of 0 and a variance of 1.
Log-normal Distribution
$\mathcal{D}\sim\text{Lognormal}(0,1)$
where the log-normal distribution is parameterized by a mean of 0 and a standard deviation of 1 for the underlying normal distribution, this too is standardized before applying the scaling.
Uniform Distribution
$\mathcal{D}\sim\mathcal{U}(-1,1)$
representing a uniform distribution between -1 and 1.
For each distribution, the synthetic returns $r_{\text{synthetic}}$ are scaled by the randomly sampled mean $\ddot{\mu}\sim\text{Uniform}(-0.001,0.001)$ and standard deviation $\ddot{\sigma}\sim\text{Uniform}(0.01,0.03)$. This scaling ensures that the synthetic datasets mimic realistic market conditions across a variety of asset types.
We generate 10 synthetic datasets per distribution, each consisting of 1000 samples, to assess the model’s robustness under these simulated conditions. Additionally, we generate $g$ synthetic datasets for the market data, where $g$ represents the number of market features. One feature is designed to have a correlation of approximately 0.7 with the synthetic returns to simulate market co-movement For more details see Appendix E.
3.4.1 Hyperparameter Optimization
The model’s hyperparameters are optimized using Bayesian optimization, with separate sets of hyperparameters for the asset-specific model (Stage 1) and the market data model (Stage 2). Each set of hyperparameters is sampled independently. The common hyperparameters and the specific hyperparameters for each stage are detailed in Table 4.
Table 4: Hyperparameters for Asset-Specific and Market Data Models
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size $b\in\{32,64,128,256,512,1024,2048\}$
Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
Normalization window $w\in[5,250]$
Dropout rate $D\in[0.0,0.9]$
L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
Use layer normalization Boolean: $\{1,0\}$
LSTM layers $l_{\text{LSTM, asset}}\in[1,5]$ $l_{\text{LSTM, market}}\in[1,5]$
LSTM units $u_{\text{LSTM, asset}}\in\{16,\cdots,256\}$ $u_{\text{LSTM, market}}\in\{16,\cdots,256\}$
Dense layers $l_{\text{dense, asset}}\in[1,5]$ $l_{\text{dense, market}}\in[1,5]$
Dense units per layer $u_{\text{dense, asset}}\in\{16,\cdots,256\}$ $u_{\text{dense, market}}\in\{16,\cdots,256\}$
[…]
3.5 Experimental Setup
Our proposed model -which will be referred to as qLSTM- is evaluated and compared against the original model architecture [2] -which will be called qDense-. It should be noted that the comparison is not entirely fair as the original model was trained under different circumstances and utilized different data. As the qDense model has a fixed forecast window of 22 days, the qLSTM model will only utilize a 22-day-ahead window as well. This means that on the test data, the assets are iterated sequentially taking 22 entries at a time, and then the loss of the 22-day-quantiles of both models is compared using the loss function described in the next section. To act as a baseline for both models, we compare them to a simple linear quantile regression (LQR) model  [^1] The LQR model is trained and tested using the same data as qDense [22].
3.5.1 Training Procedure
The model is trained using the Adam optimizer with a fixed number of 100 epochs. Early stopping is employed with a patience parameter of 10, meaning that if the validation loss does not improve for 10 consecutive epochs, training is halted. The weights that generated the lowest loss on the validation data are chosen.
3.6 Evaluation Metric
The primary evaluation metric is the quantile loss function, which has been slightly modified to disregard the volatility normalization as well as using only a single return for the evaluation. The quantile loss for a given asset is defined as
$L_{\mathcal{T}}=\frac{1}{KN}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{T}\rho_{\tau}\left(r_{t}-\hat{Q}_{r_{t}}(\tau)\right),$
where $\mathcal{T}$ is the set of quantiles used for prediction, and $K=|\mathcal{T}|$ is the number of quantiles, $T$ is the number of time steps in the test data for the given asset, $r_{t}$ is the raw return of the asset at time step $t$, and $\hat{Q}_{r_{t}}(\tau)$ is the predicted quantile of the raw return of the asset for quantile level $\tau$ at time step $t$. $\rho_{\tau}(\xi)$ is the quantile loss function, defined as:
$\rho_{\tau}(\xi)=\begin{cases}\tau\cdot\xi,&\text{if }\xi\geq 0\\ (\tau-1)\cdot\xi,&\text{if }\xi<0\end{cases}$
with $\xi=r_{i,t,j}-\hat{Q}_{r_{i,t,j}}(\tau)$ representing the residual (error) between the true return and the predicted quantile.
4 Results & Discussion
4.1 Hyper parameter Optimization
Table 5 shows the optimal hyperparameters found in the 100 trials on market data, the final model has a total of $228.390$ trainable parameters. The optimal hyperparameters of the qDense model can be found in Appendix C.
Table 5: Final Hyperparameters for the qLSTM model, rounded down to the 4th decimal place
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 256
Learning rate 0.0004
Normalization window 34
Raw LSTM layers 4 1
LSTM units 64 128
Dense layers 1 1
Dense units per layer 128 64
Dropout rate 0.1023
Market activation None
Hidden activation Leaky ReLU
Use layer normalization Yes (1)
L1 regularization 0.0009
L2 regularization 0.0010
4.2 Model Performance Evaluation
Table 6: Comparison of qDense, qLSTM, and LQR across various asset types. Percentage differences are calculated using mean percentage difference.
LQR qDense qLSTM qDense - qLSTM LQR - qLSTM
Commodities 1.4850 0.6226 0.3319 0.2908 (60.93%) 1.1532 (126.94%)
Crypto-currencies 2.5176 1.3605 0.9292 0.4313 (37.68%) 1.5884 (92.17%)
S&P 500 1.4832 0.5969 0.3284 0.2684 (58.02%) 1.1547 (127.48%)
Nikkei 225 1.4805 0.6187 0.3298 0.2889 (60.92%) 1.1507 (127.13%)
Euro Stoxx 50 1.4803 0.5983 0.3326 0.2656 (57.07%) 1.1477 (126.61%)
Currency Pairs 1.2086 0.1597 0.1156 0.0441 (32.05%) 1.0930 (165.09%)
Normal Synthetic 0.2400 0.7003 0.3549 0.3454 (65.46%) -0.1149 (-38.62%)
Log Normal Synthetic 0.4163 0.7075 0.3713 0.3362 (62.33%) 0.0450 (11.43%)
Gamma Synthetic 0.5124 0.7392 0.3805 0.3586 (64.06%) 0.1318 (29.53%)
Uniform Synthetic 0.2188 0.5683 0.2993 0.2691 (62.03%) -0.0805 (-31.06%)
Total Market Mean 1.6092 0.6594 0.3946 0.2649 (50.26%) 1.2146 (121.23%)
Total Synthetic Mean 0.3469 0.6788 0.3515 0.3273 (63.54%) -0.0046 (-1.32%)
Total Mean 1.1043 0.6672 0.3773 0.2898 (55.50%) 0.7269 (98.13%)
The results of the model evaluation across various assets and synthetic datasets, measured in quantile loss, are shown in Table LABEL:tab:final-results. Both the qLSTM and qDense models outperform the Linear Quantile Regression (LQR) model on market data. However, LQR manages to outperform both models on some of the synthetic data, as illustrated in Table LABEL:tab:final-results. The qLSTM model consistently outperforms the qDense model, with an average performance improvement of 50% on real-world datasets, 64% on synthetic datasets, and 55% overall.
Figure 1: The PDFs generated by qLSTM and qDense the first 100 out-of-sample observations of BTC-USD
On real-world data, the qLSTM model demonstrates a significant performance boost. For instance, when applied to S&P 500 stocks, the qLSTM model achieves a quantile loss of 0.3284, compared to 0.5969 for the qDense model and 1.4805 for LQR—representing a 58% and 127% improvement, respectively. Similarly, for cryptocurrencies, the qLSTM model reduces quantile loss by 37% compared to qDense and 92% compared to LQR, showcasing its robustness in handling volatile and unpredictable assets. Figure 1 shows the distribution of quantile losses for qLSTM and qDense over time for the real-world data. The qDense distribution undergoes greater shape variation and appears more symmetrical, while the qLSTM model captures more complexity with less variation in shape, indicating its ability to model intricate distribution patterns. This suggests that qLSTM captures dynamic distributions more effectively, whereas qDense tends to produce distributions closer to normality. This is further illustrated in Figure 2, where a single slice of the distributions shows qDense’s smoother, nearly normal distribution, while qLSTM displays a more intricate and asymmetrical shape, which is better aligned with real data.
Figure 2: The PDFs generated by qLSTM and qDense on a randomly chosen out-of-sample observation of BTC-USD
When evaluated on synthetic data, the qLSTM model shows greater improvement over qDense, but still falls short of LQR on average. qLSTM outperforms LQR on log-normal and gamma distributions, improving quantile loss by 11% and 30%, respectively, demonstrating its ability to capture asymmetrical distributions effectively. However, on the normal and uniform distributions, which are largely symmetrical, LQR outperforms qLSTM by 39% and 31%, respectively. This is somewhat expected, given the nice, symmetric behavior of these distributions. Figure 3 shows the first 100 samples of synthetically generated normally distributed data, where neither distribution changes significantly in shape, with LQR showing no notable variation, as expected for i.i.d. data. This reflects how well the models capture the structure of synthetic data. Although LQR performs better on symmetrical distributions, qLSTM still generates somewhat asymmetrical distributions that approach normality. Figure 4 presents a single slice of the distributions produced by LQR and qLSTM, highlighting qLSTM’s simpler distribution in comparison to real-world data, though still exhibiting slight asymmetry. Interestingly, LQR shows a more complex shape in this case, despite achieving lower quantile loss than qLSTM on this dataset.
Figure 3: The PDFs generated by qLSTM and LQR on the first 100 synthetic normally distributed data
Figure 4: The PDFs generated by qLSTM and LQR on a randomly chosen observation of synthetic normally distributed data
4.3 Limitations and Challenges
While our approach yielded promising results, several limitations persist. One significant limitation is the inability to fairly compare our proposed approach with that of the authors in [2]. This issue arises primarily due to the lack of access to the original dataset used in their study and the absence of publicly available code for their implementation. Consequently, we were unable to replicate their exact methodology, which inherently limits the comparison between qLSTM approach and their dense model.
Furthermore, the two approaches differ fundamentally in how they process data. The qLSTM model is specifically designed to handle sequence data, allowing it to capture patterns across variable-length sequences and predict multiple quantiles for future time steps simultaneously. In contrast, the dense model used in [2] operates on fixed-length windows and focuses on single-step quantile predictions. This introduces a degree of inconsistency in the comparison, as the qLSTM’s ability to process sequential information may provide it with a natural advantage in capturing long-term dependencies in asset returns.
Another challenge emerged when comparing model performance on synthetic data. Due to differences in code constraints and hyperparameter configurations, the two models were not exposed to the same number of observations. Specifically, the qLSTM model used a normalization window of 34 days, while the qDense model operated on a longer window of 123 days. This discrepancy means that the qLSTM model was trained on a larger number of observations, increasing its potential for failure. However, despite being exposed to more opportunities for failure, the qLSTM model still outperformed the qDense model.
5 Conclusion
In this paper, we presented a novel model for predicting the distribution of asset returns using a quantile-based approach, extending prior research. By leveraging Long Short-Term Memory (LSTM) networks, we modeled the distribution of returns more dynamically and generalized these predictions across multiple asset classes, including synthetic data. Our approach diverged from previous models by incorporating asset-neutral features and utilizing market data to further refine the predictions.
The results demonstrated significant improvements over a baseline linear quantile regression model, outperforming it by 98% across all datasets. Additionally, the LSTM approach outperformed a dense network model by over 50% on both real-world and synthetic data. This performance boost underscores the strength of our approach in capturing the complexities of financial return distributions, particularly for high-volatility assets like cryptocurrencies. Furthermore, the model’s ability to maintain strong performance across diverse synthetic datasets highlights its generalizability and flexibility.
For future work, expanding the dataset to include more assets and asset classes—such as bonds, options, or credit default swaps—could further test the model’s generalizability. Additionally, exploring different time intervals, such as hourly, minute, or even second-level data, could provide deeper insights, as this study exclusively utilized daily intervals. Testing the model on illiquid assets could further validate its robustness in different market conditions. In addition, incorporating alternative data sources, such as social media sentiment, news feeds, or satellite imagery, might further enhance model accuracy. Finally, future research could focus on optimizing the LSTM-based model for real-world applications, such as integrating it into real-time trading systems or risk management frameworks.
Statements and Declarations
The authors declare that there are no competing interests related to this work.
Appendix A Financial Assets
The table below provides a comprehensive list of financial assets, including their ticker symbols, associated indexes, start dates, annualized mean returns, and standard deviations of returns, rounded to two decimal places where applicable.
Table 7: List of assets with their tickers, index, start date, annualized mean return, and standard deviation.
Name Ticker Index Start Date Mean (%) Std (%)
Apple Inc. AAPL S&P 500 2000-02-03 1.46 40.43
Microsoft Corporation MSFT S&P 500 2000-02-03 0.66 30.49
Amazon.com, Inc. AMZN S&P 500 2000-02-03 1.00 49.22
Alphabet Inc. (Class A) GOOGL S&P 500 2004-09-21 1.26 30.49
The Walt Disney Company DIS S&P 500 2000-02-03 0.30 30.42
Johnson & Johnson JNJ S&P 500 2000-02-03 0.51 19.16
Visa Inc. V S&P 500 2008-04-21 1.13 29.07
NVIDIA Corporation NVDA S&P 500 2000-02-03 1.72 59.48
Mastercard Inc. MA S&P 500 2006-06-27 1.65 33.40
The Coca-Cola Company KO S&P 500 2000-02-03 0.37 20.58
TotalEnergies SE TOTF.PA Euro Stoxx 50 2000-02-02 0.43 26.51
Volkswagen AG VOW3.DE Euro Stoxx 50 2000-02-02 0.62 37.29
L’Oréal SA OR.PA Euro Stoxx 50 2000-02-02 0.59 25.12
BNP Paribas SA BNP.PA Euro Stoxx 50 2000-02-02 0.42 36.88
Daimler AG DAI.DE Euro Stoxx 50 2000-02-02 0.57 32.36
Siemens AG SIE.DE Euro Stoxx 50 2000-02-02 0.27 35.95
Airbus SE AIR.PA Euro Stoxx 50 2001-10-03 0.83 37.68
Muenchener Rueckversicherungs AG MUV2.DE Euro Stoxx 50 2000-02-02 0.31 30.00
Intesa Sanpaolo S.p.A. INTC.MI Euro Stoxx 50 2000-02-02 0.35 39.94
BASF SE BAS.DE Euro Stoxx 50 2000-02-02 0.47 29.07
Toyota Motor Corporation 7203.T Nikkei 225 2000-02-02 0.41 28.49
Sony Group Corporation 6758.T Nikkei 225 2000-02-03 0.13 36.51
SoftBank Group Corp. 9984.T Nikkei 225 2000-02-03 0.06 51.28
Nintendo Co., Ltd. 7974.T Nikkei 225 2001-02-05 0.51 38.81
Mitsubishi UFJ Financial Group, Inc. 8306.T Nikkei 225 2005-10-31 0.13 33.08
Tokyo Electron Limited 8035.T Nikkei 225 2000-02-03 0.53 42.49
Mitsubishi Corporation 8058.T Nikkei 225 2000-02-03 0.71 33.12
HOYA Corporation 7741.T Nikkei 225 2001-01-31 0.68 32.51
Fast Retailing Co., Ltd. 9983.T Nikkei 225 2000-02-03 0.72 42.37
Omron Corporation 6645.T Nikkei 225 2001-02-05 0.41 35.85
British Pound to US Dollar GBPUSD=X Currency Pairs 2003-12-31 -0.10 9.54
US Dollar to Japanese Yen USDJPY=X Currency Pairs 2000-02-02 0.07 11.34
Australian Dollar to US Dollar AUDUSD=X Currency Pairs 2006-06-15 -0.03 59.98
Euro to British Pound EURGBP=X Currency Pairs 2000-02-02 0.09 8.72
Euro to Japanese Yen EURJPY=X Currency Pairs 2003-02-24 0.06 11.48
Euro to Swiss Franc EURCHF=X Currency Pairs 2003-02-24 -0.13 7.89
Euro to Australian Dollar EURAUD=X Currency Pairs 2003-12-31 -0.01 10.39
US Dollar to Indian Rupee USDINR=X Currency Pairs 2003-12-31 0.18 7.76
Swiss Franc to Japanese Yen CHFJPY=X Currency Pairs 2003-12-31 0.20 11.46
US Dollar to Hong Kong Dollar USDHKD=X Currency Pairs 2001-08-15 0.00 1.07
Copper Futures HG=F Commodities 2000-10-02 0.39 26.89
Wheat Futures ZW=F Commodities 2000-08-16 0.26 32.82
Coffee Futures KC=F Commodities 2000-02-03 0.14 34.01
Cocoa Futures CC=F Commodities 2000-02-03 0.45 30.28
Cotton Futures CT=F Commodities 2000-02-03 0.09 30.29
Live Cattle Futures LE=F Commodities 2001-04-02 0.23 18.43
Platinum Futures PL=F Commodities 2000-02-04 0.22 31.50
Palladium Futures PA=F Commodities 2000-02-04 0.21 38.62
Corn Futures ZC=F Commodities 2000-08-16 0.27 29.14
Orange Juice Futures JO=F Commodities 2001-10-17 0.37 33.20
Bitcoin BTC-USD Cryptocurrencies 2014-10-09 2.25 59.09
Ethereum ETH-USD Cryptocurrencies 2017-12-01 1.16 76.15
Ripple XRP-USD Cryptocurrencies 2017-12-01 0.64 96.11
Cardano ADA-USD Cryptocurrencies 2017-12-01 1.14 94.15
Dogecoin DOGE-USD Cryptocurrencies 2017-12-01 2.69 115.10
Binance Coin BNB-USD Cryptocurrencies 2017-12-01 3.61 86.33
NEAR Protocol NEAR-USD Cryptocurrencies 2020-11-05 2.65 109.42
Bitcoin Cash BCH-USD Cryptocurrencies 2017-12-01 -1.20 93.81
Litecoin LTC-USD Cryptocurrencies 2014-10-09 1.38 85.60
Solana SOL-USD Cryptocurrencies 2020-05-02 5.85 114.08
Appendix B Market Data
The table below lists the market data used in this study, including tickers, a brief representation, start dates, annualized mean returns, and standard deviations. This data spans a variety of indices, futures, and financial instruments, providing a broad view of market performance and trends over time.
Table 8: Market data with tickers, representation, start date, annualized mean, and standard deviation.
Name Ticker Representation Start Date Mean (%) Std (%)
S&P 500 Index ^GSPC U.S. stock market performance 2000-02-03 0.50 21.05
NASDAQ Index ^IXIC Tech-focused stock performance 2000-02-03 0.75 23.29
Dow Jones Index ^DJI Large U.S. company performance 2000-02-03 0.46 19.93
Russell 2000 Index ^RUT Small-cap U.S. stock performance 2000-02-03 0.38 26.65
Euro Stoxx 50 Index ^STOXX50E Top 50 European company performance 2000-02-03 0.03 23.52
Nikkei 225 Index ^N225 Japanese stock market performance 2000-02-03 0.27 24.07
Crude Oil Futures CL=F Global economic activity indicator 2000-02-03 0.28 46.21
Gold Futures GC=F Inflation hedge and safe haven 2000-02-03 0.47 18.39
Silver Futures SI=F Industrial demand and inflation hedge 2000-02-03 0.24 34.22
VIX Index ^VIX Market volatility (fear index) 2000-02-03 -0.07 125.89
10 Year Treasury Yield ^TNX Long-term interest rates and growth 2000-02-03 -0.08 47.40
3 Month Treasury Yield ^IRX Short-term interest rates and liquidity 2000-02-03 -0.28 478.84
US Dollar Index DX-Y.NYB U.S. dollar strength against a basket of currencies 2000-02-03 0.08 8.12
Corporate Bond Index AGG U.S. investment-grade bond market performance 2000-02-03 0.20 5.75
Appendix C qDense Training Details & Results
The following tables summarize the hyperparameters used during the training of the qDense model. The first table presents the ranges explored during the hyperparameter tuning process, and the second table lists the final hyperparameters used in the model, rounded down to the fourth decimal place.
Table 9: Hyperparameters explored for qDense Asset-Specific and Market Data Models
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size $b\in\{32,64,128,256,512,1024,2048\}$
Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
Normalization window $w\in[5,250]$
Dropout rate $D\in[0.0,0.9]$
L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
Use layer normalization Boolean: $\{1,0\}$
Raw hidden layers $l_{\text{raw, asset}}\in[1,5]$ $l_{\text{raw, market}}\in[1,5]$
Hidden layer units $u_{\text{raw, asset}}\in\{16,\cdots,512\}$ $u_{\text{raw, market}}\in\{16,\cdots,512\}$
[…]
Table 10: Final Hyperparameters for the qDense model, rounded down to the 4th decimal place
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 2048
Learning rate 0.0007
Normalization window 123
Raw hidden layers 2 4
Hidden layer units 64, 512 64, 512, 16, 32
Dropout rate 0.1780
Market activation Sigmoid
Hidden activation ELU
Use layer normalization Yes (1)
L1 regularization 0.0007
L2 regularization 0.0004
Appendix D Kernel Density Estimation
Kernel Density Estimation (KDE) is a non-parametric method used to estimate the probability density function (PDF) of a random variable. In this study, KDE was used to convert predicted quantiles into a smooth probability density function. The advantage of KDE is that it allows for flexible modeling of the underlying data distribution without assuming any particular parametric form, which is useful when working with financial return distributions that may not follow standard distributions like normal or log-normal.
The quantiles predicted by the model are transformed into a smooth PDF using KDE, ensuring that the final distribution is continuous and appropriately represents the data. This process helps provide more accurate density estimates, which are crucial for downstream tasks such as risk estimation and derivative pricing.
Below is the Python code used for the KDE conversion in this work.
1import numpy as np
2import matplotlib.pyplot as plt
3from sklearn.neighbors import KernelDensity
4
5def quantiles_to_pdf_kde(quantiles, grid_points=200, bandwidth=0.01):
6 """
7 Convert quantiles to a probability density function (PDF) using Kernel Density Estimation (KDE).
8
9 Parameters:
10 - quantiles: Array of quantile values.
11 - grid_points: Number of grid points for a denser representation.
12 - bandwidth: Bandwidth for the KDE (controls the smoothness of the density).
13
14 Returns:
15 - pdf_grid: Probability density function values on the dense grid.
16 - cdf_grid: Cumulative distribution function values on the dense grid (obtained by integrating the PDF).
17 - dense_grid: The dense grid of quantiles corresponding to the PDF and CDF values.
18 """
19
20 # Create a denser grid between the quantiles
21 dense_grid = np.linspace(min(quantiles), max(quantiles), grid_points).reshape(-1, 1)
22
23 # Fit KDE to the quantile values
24 kde = KernelDensity(kernel=’gaussian’, bandwidth=bandwidth).fit(quantiles.reshape(-1, 1))
25
26 # Compute log density and convert to regular density
27 log_dens = kde.score_samples(dense_grid)
28 pdf_grid = np.exp(log_dens)
29
30 # Approximate the CDF by cumulatively summing the PDF
31 cdf_grid = np.cumsum(pdf_grid)
32 cdf_grid /= cdf_grid[-1] # Normalize to [0, 1]
33
34 return pdf_grid, cdf_grid, dense_grid
Listing 1: Quantiles to PDF using KDE
Appendix E Synthetic Data
In this study, synthetic data was generated across four different distributions: Normal, Log-Normal, Gamma, and Uniform. For each distribution, multiple assets were created with specific $\mu$ (mean) and $\sigma$ (standard deviation) values. These synthetic datasets were designed to match a target correlation with real-world market data, ensuring that the generated data maintained realistic statistical properties. The table below summarizes the $\mu$, $\sigma$, and correlation values for the synthetic assets across the various distributions.
Table 11: Generated $\mu$ and $\sigma$ values and correlation with market data across different test datasets.
Distribution Type Asset Generated $\mu$, $\sigma$ Correlation with Market Data
Normal Asset 1 $\mu=-0.000663$, $\sigma=0.022852$ 0.7047
Asset 2 $\mu=-0.000944$, $\sigma=0.027103$ 0.6787
Asset 3 $\mu=0.000427$, $\sigma=0.010072$ 0.7149
Asset 4 $\mu=-0.000551$, $\sigma=0.020746$ 0.6908
Asset 5 $\mu=-0.000796$, $\sigma=0.023492$ 0.7261
Asset 6 $\mu=-0.000633$, $\sigma=0.018013$ 0.7077
Asset 7 $\mu=0.000044$, $\sigma=0.028276$ 0.6712
Asset 8 $\mu=-0.000502$, $\sigma=0.017238$ 0.6884
Asset 9 $\mu=0.000452$, $\sigma=0.021794$ 0.7028
Asset 10 $\mu=0.000647$, $\sigma=0.012101$ 0.6820
Log Normal Asset 1 $\mu=0.000922$, $\sigma=0.015048$ 0.6900
Asset 2 $\mu=0.000839$, $\sigma=0.029755$ 0.7013
Asset 3 $\mu=-0.000659$, $\sigma=0.025729$ 0.6859
Asset 4 $\mu=-0.000637$, $\sigma=0.028419$ 0.6892
Asset 5 $\mu=0.000824$, $\sigma=0.029507$ 0.7040
Asset 6 $\mu=-0.000740$, $\sigma=0.010353$ 0.7052
Asset 7 $\mu=-0.000874$, $\sigma=0.025497$ 0.6918
Asset 8 $\mu=0.000052$, $\sigma=0.016105$ 0.6987
Asset 9 $\mu=-0.000106$, $\sigma=0.022525$ 0.7048
Asset 10 $\mu=0.000704$, $\sigma=0.024870$ 0.6874
Gamma Asset 1 $\mu=-0.000725$, $\sigma=0.013788$ 0.6954
Asset 2 $\mu=0.000544$, $\sigma=0.018391$ 0.7030
Asset 3 $\mu=0.000613$, $\sigma=0.023223$ 0.7012
Asset 4 $\mu=0.000029$, $\sigma=0.015418$ 0.7063
Asset 5 $\mu=-0.000095$, $\sigma=0.026113$ 0.7214
Asset 6 $\mu=-0.000846$, $\sigma=0.010134$ 0.7045
Asset 7 $\mu=0.000724$, $\sigma=0.028490$ 0.6717
Asset 8 $\mu=0.000584$, $\sigma=0.026327$ 0.7016
Asset 9 $\mu=0.000503$, $\sigma=0.029417$ 0.7100
Asset 10 $\mu=-0.000688$, $\sigma=0.029474$ 0.6916
Uniform Asset 1 $\mu=-0.000603$, $\sigma=0.014227$ 0.8650
Asset 2 $\mu=-0.000417$, $\sigma=0.016919$ 0.8603
Asset 3 $\mu=-0.000024$, $\sigma=0.013627$ 0.8617
Asset 4 $\mu=-0.000554$, $\sigma=0.026960$ 0.8579
Asset 5 $\mu=0.000231$, $\sigma=0.026206$ 0.8597
Asset 6 $\mu=-0.000763$, $\sigma=0.028292$ 0.8590
Asset 7 $\mu=0.000122$, $\sigma=0.014288$ 0.8603
Asset 8 $\mu=0.000404$, $\sigma=0.017516$ 0.8587
Asset 9 $\mu=-0.000082$, $\sigma=0.026868$ 0.8610
Asset 10 $\mu=-0.000876$, $\sigma=0.012219$ 0.8545
In addition to the statistical properties shown in the table, the synthetic data was generated using the following Python function. The function allows for the generation of log returns that are correlated with real market data by controlling the mean ($\mu$) and standard deviation ($\sigma$) of the generated data.
1import numpy as np
2import pandas as pd
3from scipy.stats import gamma, lognorm, uniform
4
5def generate_synthetic_data_with_target_correlation(market_data, target_correlation, mu, sigma, distribution=’normal’):
6 """
7 Generate a synthetic dataset with returns that have a target correlation to the market_data.
8
9 Parameters:
10 - market_data: numpy array of log returns from the market data
11 - target_correlation: the target correlation for synthetic data’s log returns to the market data
12 - mu: mean of the synthetic dataset
13 - sigma: standard deviation of the synthetic dataset
14 - distribution: the distribution to sample synthetic data from (’normal’, ’gamma’, ’lognormal’, ’uniform’)
15
16 Returns:
17 - df: DataFrame with synthetic prices
18 """
19
20 num_periods = len(market_data)
21
22 # Step 1: Generate independent noise based on the specified distribution
23 if distribution == ’normal’:
24 noise = np.random.normal(loc=0, scale=1, size=num_periods) # Standard normal noise
25 elif distribution == ’gamma’:
26 shape = 2 # Shape parameter for the gamma distribution
27 noise = gamma.rvs(a=shape, scale=1, size=num_periods)
28 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
29 elif distribution == ’lognormal’:
30 noise = lognorm.rvs(s=1, scale=np.exp(0), size=num_periods)
31 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
32 elif distribution == ’uniform’:
33 noise = uniform.rvs(loc=-1, scale=2, size=num_periods) # Uniform distribution [-1, 1]
34 else:
35 raise ValueError("Unsupported distribution. Choose from ’normal’, ’gamma’, ’lognormal’, or ’uniform’.")
36
37 # Step 2: Standardize market data
38 market_std = (market_data - np.mean(market_data)) / np.std(market_data)
39
40 # Step 3: Create a linear combination of market data and noise for correlated returns
41 synthetic_log_returns = target_correlation * market_std + np.sqrt(1 - target_correlation**2) * noise
42
43 # Step 4: Adjust the synthetic returns to have the desired mean and standard deviation
44 synthetic_log_returns = synthetic_log_returns * sigma + mu
45
46 # Step 5: Convert synthetic log returns to price data
47 initial_price = np.random.uniform(5, 1000) # Random initial price
48 synthetic_prices = initial_price * np.exp(np.cumsum(synthetic_log_returns))
49
50 # Step 6: Create a DataFrame with synthetic price data
51 df = pd.DataFrame(synthetic_prices, columns=[’Adj Close’])
52
53 # Add High, Low, Volume columns to allow all functions to work
54 df[’High’] = df[’Adj Close’] * (1 + np.random.normal(0, 0.01, size=len(df)))
55 df[’Low’] = df[’Adj Close’] * (1 - np.random.normal(0, 0.01, size=len(df)))
56 df[’Volume’] = 0
57
58 df = generate_technical_features(df)
59
60 return df
Listing 2: Synthetic Data Generation with Target Correlation
The function generates synthetic log returns that are then converted to price data, while ensuring a target correlation with the real market data. This allows for a more controlled and realistic evaluation of the model’s performance on different types of financial data.
Appendix F Quantiles
The quantiles used in this study are:
[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03,[ 0.00005 , 0.00025 , 0.00075 , 0.00125 , 0.00175 , 0.0025 , 0.005 , 0.01 , 0.015 , 0.02 , 0.03 ,
$0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7,$
0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995 0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995]0.75 , 0.8 , 0.85 , 0.9 , 0.95 , 0.98 , 0.99 , 0.995 , 0.9975 , 0.99925 , 0.99975 , 0.99995 ]

### v2

@@
-Generalized Distribution Prediction for Asset Returns
+Generalized Distribution Estimation for Asset Returns
 Abstract
-We present a novel approach for predicting the distribution of asset returns using a quantile-based method with Long Short-Term Memory (LSTM) networks.
-Our model is designed in two stages: the first focuses on predicting the quantiles of normalized asset returns using asset-specific features, while the second stage incorporates market data to adjust these predictions for broader economic conditions.
-This results in a generalized model that can be applied across various asset classes, including commodities, cryptocurrencies, as well as synthetic datasets.
-The predicted quantiles are then converted into full probability distributions through kernel density estimation, allowing for more precise return distribution predictions and inferencing.
-The LSTM model significantly outperforms a linear quantile regression baseline by 98% and a dense neural network model by over 50%, showcasing its ability to capture complex patterns in financial return distributions across both synthetic and real-world data.
-By using exclusively asset-class-neutral features, our model achieves robust, generalizable results.
+Accurately modeling the distribution of asset returns enables better risk management, portfolio optimization, and financial decision-making especially in uncertain and volatile markets.
+This paper presents a novel approach to estimating the distribution of asset log returns using quantile regression combined with smoothed density estimation.
+Our proposed model, qLSTM, leverages asset-neutral features to deliver robust and generalizable predictions across diverse asset classes, outperforming baseline and dense models in quantile loss metrics for real-world data.
+A hybrid model, averaging Gaussian and model-predicted quantiles, qHybrid, further improves performance on synthetic datasets with skewed and heavy-tailed distributions.
+Evaluation using metrics like Wasserstein distance, CRPS, and VaR demonstrates the models’ effectiveness in capturing tail risks and broader distributional characteristics.
 keywords:
-Machine Learning, Distribution Prediction, Asset Return Forecasting, Long Short-Term Memory (LSTM) Networks
+Distribution Estimation , Qauntile Regression , LSTM , Tail Risk , Financial Modeling , Machine Learning
+journal: Applied Soft Computing
+[…]
+[…]
 1 Introduction
-Every day, an almost uncountable number of transactions take place in capital markets, with each transaction generally aimed at generating a positive return on investment.
-If one could predict future financial asset prices accurately, the potential for gains would be nearly limitless.
-Even achieving a prediction accuracy slightly above 50% could potentially turn someone into a billionaire [35].
-However, predicting returns is an extraordinarily difficult problem.
-Nobel laureate Eugene Fama famously argued that future returns are inherently random and, therefore, practically impossible to predict [11].
-This difficulty is further supported by the fact that many funds consistently fail to outperform the S&P 500, a common benchmark for returns [13, 19].
-In finance, logarithmic asset returns are often assumed to follow a normal distribution, a famous example is the Black-Scholes model which is used to find the fair value of option derivatives [3].
+In the financial world, accurate modeling of asset price movements and their associated risks is a cornerstone of decision-making for investors, portfolio managers, and risk analysts [1].
+This modeling forms the foundation of derivative pricing, risk management, and portfolio optimization.
+One widely used approach assumes that asset price changes over time can be modeled as a stochastic process, with the distribution of returns playing a central role in these models [2].
+Machine learning has transformed several practices and processes withing the finance domain by enabling models that capture the complex, nonlinear relationships inherent in market data [3].
+Applications range from price and return prediction to risk management and portfolio optimization, leveraging methods such as ARIMA, LSTM, and CNNs [4, 5, 6].
+However, much of this research focuses on point predictions rather than estimating the full return distribution, which is essential for understanding uncertainty and tail risks.
+Traditionally, logarithmic asset returns are often assumed to follow a normal distribution; a famous example is the Black-Scholes model, which is used to find the fair value of option derivatives [2].
 However, in practice, this assumption frequently does not hold [7].
@@
 Consequently, models that rely on normality assumptions can lead to suboptimal or even incorrect results [8].
-To address this issue, some approaches generalize assumptions or adopt alternative distributions, such as the Student’s T distribution, which accommodates skewness, kurtosis, and heavy tails [1].
+To address this issue, some approaches generalize assumptions or adopt alternative distributions, such as the Student’s $t$- distribution, which accommodates skewness, kurtosis, and heavy tails [9].
+Furthermore, distributional approaches in machine learning, show recent advances in conditional modeling and quantile regression, address this gap by providing a stronger method for distribution estimation [10, 11].
 More recently, a new approach has emerged, namely to model the distribution directly [12].
@@
 A more accurate model of the return distribution could improve derivative pricing, inform risk management, and enhance portfolio optimization.
-In this paper, we propose a model that uses a distributional approach to predict asset returns.
-Our work extends the model presented in [2], which predicts the distribution of stock returns.
-Instead of predicting the return itself, our model predicts the quantiles of the return distribution.
-Although quantile regression is not a novel concept, we convert these quantiles into a full distribution using kernel density estimation.
+Although the proposed method shows promising results, it only considers one day’s worth of data and has a fixed look ahead period of 22 days.
+Furthermore, the method relies heavily on derived features which are exclusive to stocks, meaning the method does not generalize to other asset classes.
+In this paper, we propose a model that estimates the distribution of returns given current observations.
+Our model predicts the quantiles of the log return distribution over $n$ days given $n$ days worth of data, thus extending previous work[12].
+Although quantile regression is not a novel concept, we convert these quantiles into a full distribution using a smoothed density estimation function.
 This enables a more precise estimation of the return distribution over a given time period.
-The model is trained on a dataset of asset returns and general market variables, ensuring that each feature is asset-class neutral, making the approach applicable across different asset classes.
-This paper contributes to the literature by creating a more generalizable distribution prediction model, utilizing only asset-neutral features and allowing for the consideration of multiple return types.
-The results are compared to a baseline linear quantile regression model and the original approach in [2].
-We find that our generalized approach outperforms the baseline by 121% on real-world data, while slightly underperforming by 1% on synthetic data.
-Compared to the original approach, our model outperforms on both real-world and synthetic data by 50% and 64%, respectively.
+The model is trained on a dataset of asset returns and general market variables, ensuring that each characteristic is asset-class neutral, which makes the approach applicable across different asset classes.
+This paper contributes to the literature in the following ways:
+  • 1.
+    Proposes a generalizable distribution estimation model that:
+    • (a)
+      Uses only asset-neutral features.
+    • (b)
+      Allows the consideration of multiple periods of return.
+  • 2.
+    Proposes a hybrid model that utilizes standard Gaussian quantiles in combination with the generalized model.
+  • 3.
+    Introduces an experimental framework for enhanced reproducibility in the field by using only data which is is publicly available.
+The results are compared in two ways: quantile loss is used to evaluate the performance of our approaches against a baseline linear quantile regression model and the original dense approach in [12].
+We find that our proposed qLSTM model achieves the lowest overall quantile loss for real-world data, demonstrating its robustness and versatility across diverse asset classes, including cryptocurrencies, commodities, and equity indices.
+The hybrid model, qHybrid, performs particularly well on synthetic data, especially for distributions with heavy tails or skewness, highlighting its flexibility.
+Additionally, we evaluate the models using standard metrics such as the Wasserstein distance and the Continuous Ranked Probability Score (CRPS).
+These metrics reveal that qLSTM excels in capturing extreme tail risks, while the qHybrid model achieves better calibration for broader distributions.
+The rest of this paper is structured as follows.
+In Section 2, we provide an overview of the existing literature on machine learning in finance, focusing on the state of machine learning in finance, distributional modeling, and quantile regression methods.
+Section 3 outlines our methodology, including the problem formulation, model architecture, and training procedure.
+The data sources, including real-world and synthetic datasets, are described in detail, along with feature engineering and hyperparameter optimization.
+Section 4 presents the results of our experiments, comparing the performance of our proposed models with existing baselines using various metrics.
+Finally, Section 5 concludes the paper by summarizing our contributions and findings.
 2 Literature Review
-2.1 Asset Return Prediction
-The field of asset return prediction using machine learning and artificial intelligence has expanded rapidly in recent years.
-The sheer volume of published studies makes it increasingly difficult to conduct a thorough review of the existing literature.
-As noted by Dessain, the number of academic studies has grown at an exponential rate, creating challenges for researchers attempting to synthesize existing knowledge [8].
-He observes that conducting a comprehensive review of the literature on machine learning in finance is "becoming more and more challenging, if feasible at all," due to the rapid influx of studies [8].
+The application of machine learning to estimate phenomena within finance is well-explored, with much of the literature focusing on optimizing returns by predicting directions, prices, or returns.
+As noted by Dessain [13], the exponential growth in academic studies has made it increasingly difficult to synthesize existing knowledge within the domain.
 Similarly, Strader et al.
-[31] highlighted the difficulty in comparing studies because of the vast diversity of approaches and benchmarks used across research.
-In addition to the sheer volume of publications, the complexity and diversity of machine learning models used in asset return prediction further complicates synthesis.
-The review by Shah et al.
-underscores the variety of methods employed, including ARIMA [20, 9, 25], LSTM [28, 34, 21], CNN [5, 15, 17], and hybrid models [32, 24, 14].
-This wide range of methodologies makes it challenging to compare results across studies [29].
-This issue aligns with Dessain’s observation that conducting a comprehensive review is exceptionally difficult [8].
-Another factor that hinders a thorough and consistent literature review is the variation in data used across studies [8].
-Differences in asset classes, individual assets, time periods, and time intervals, often chosen arbitrarily, complicate meaningful comparisons between methodologies.
-To mitigate this, our paper uses publicly available data, ensuring that our results can be easily compared to other studies.
-While several studies have attempted to address these challenges [27, 18, 4], they too struggle to keep pace with the rapid rate of publication and face the same issues described above.
-These efforts are important in their own right, but the complexities of the field continue to pose significant challenges for comprehensive reviews.
-2.2 Distribution Prediction
-Distribution prediction is a relatively novel approach in the machine learning literature.
-One of the few papers to focus on this is [2], where the authors propose a model that predicts quantiles of the return distribution rather than the return itself.
-This approach aligns with the quantile regression framework discussed by Liu [23].
-The concept has been further advanced by models like the Conditional Quantile Variational Autoencoder (CQVAE), introduced by Yang et al.
-[33], which employs neural networks to model conditional quantiles.
-Together, these studies demonstrate the potential of distribution prediction and highlight its advantages over more traditional methods, which are often constrained by certain assumptions about return distributions.
-Despite these advances, there remains significant room for improvement, and this is where our work contributes.
-While [2] predicts the 22-day ahead quantiles of returns for a specific asset, our model leverages Long Short-Term Memory (LSTM) networks to predict quantiles for an $n$-day period, where $n$ corresponds to both the input sequence length and the forecast window.
-Our approach also differs in the type of input features used.
-The model in [2] incorporates 194 stock-specific characteristics along with macroeconomic variables to predict return quantiles.
-In contrast, our model employs asset-neutral features, drawing on the generalization techniques from the CQVAE model to ensure robustness and applicability across various asset classes.
+[14] and Shah et al..
+[15] emphasize the diversity of methods, such as ARIMA [4, 16], LSTM [5, 17], CNN [6, 18], and hybrid models [19, 20], as well as the inconsistency in the benchmarks, which complicates the comparisons between studies.
+To address these issues, our study uses publicly available data to ensure reproducibility and comparability.
+In the domain of distribution estimation, significant advances have been made in non-conditional density estimation [21, 10, 22].
+In contrast, this work focuses on conditional distributions allowing us to take various features from both the asset in question as well as features from the market to inform the shape of the distribution.
+Hu and Nan propose a method for estimating conditional distributions by modeling the hazard function and deriving the cumulative distribution function (CDF) via integration [11], while Feindt [23] demonstrates the utility of CDF-based methods across domains, including finance.
+However, both approaches prioritize global density estimation.
+Our work differs by directly estimating the quantile function, emphasizing financial applications, and incorporating domain-specific knowledge to capture tail risks.
+Another notable contribution is by Rasul et al..
+[10], who use normalizing flows conditioned on time series data for multistep distributional estimation, where the conditional distribution of $r_{k}$ depends not only on past observations $X$, but also on intermediate estimations $r_{1},\ldots,r_{k-1}$.
+This allows uncertainty to propagate across horizons.
+In contrast, we assume that each return on the forecast trajectory is drawn from the same distribution (i.e., $r_{1}\sim f(X),r_{2}\sim f(X)$), simplifying the computation while focusing on quantile estimation and tail risk for financial applications.
+Focusing specifically on finance, Hickling and Prangle [24] leverage extreme value theory inspired transformations within normalizing flows to model heavy-tailed distributions.
+Similarly, Barun and Liu.
+[25] utilize quantile regression to address distributional challenges in financial returns.
+Building on this foundation, our work incorporates long-short-term memory (LSTM) networks to estimate quantiles over $n$ day horizons, ensuring robustness across asset classes by utilizing asset-neutral features, in contrast to feature-heavy approaches like those in [12], which rely on stock-specific characteristics and macroeconomic variables.
+Despite progress, challenges remain in balancing model complexity, interpretability, and applicability across asset classes.
+This work addresses gaps in tail risk modeling by combining quantile-based LSTM architectures with asset-neutral features, providing robust and generalizable estimations across diverse financial datasets.
 3 Methodology
-In this study, we employ a two-stage model to predict quantiles of asset returns.
-The first stage uses only asset-specific data to predict quantiles for normalized returns.
-The second stage incorporates market data to generate a scaling factor, which is then multiplied by the output from the first stage.
-This scaling factor adjusts the normalized returns based on broader market conditions.
-To train the model, we utilize a modified quantile loss function that accommodates multiple observations, in line with the generalized approach.
-This allows for a more robust evaluation of the model’s predictive capability across different datasets.
-We validate the methodology using both real-world and synthetic data.
-The code used to implement all approaches in this section as well as to evaluate the models is publicly available at https://github.com/izzak98/dist-pred
+Our proposed methodology consists of several steps.
+First we formalize the problem which this paper tackles, then we describe our two stage model to take into account both asset specific features as well as market general features.
+It is then followed by details about the two stage quantile loss function deployed.
+Next, we describe both real and synthetic data sources used in the training and evaluation of the models.
+Last, we detail the experimental setup and the metrics used for evaluation.
+The code used to implement all the approaches in this section, as well as to evaluate the models, is publicly available at https://github.com/izzak98/dist-pred
 3.1 Problem Definition
@@
 The matrix $X$ is structured as follows
-$X=\begin{bmatrix}x_{1,1}&x_{1,2}&\cdots&x_{1,t}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,t}\\ \vdots&\vdots&\ddots&\vdots\\ x_{m,1}&x_{m,2}&\cdots&x_{m,t}\\ \end{bmatrix}$
+$X=\begin{bmatrix}x_{1,1}&x_{1,2}&x_{1,3}&\cdots&x_{1,t}\\ x_{2,1}&x_{2,2}&x_{2,3}&\cdots&x_{2,t}\\ x_{3,1}&x_{3,2}&x_{3,3}&\cdots&x_{3,t}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ x_{m,1}&x_{m,2}&x_{m,3}&\cdots&x_{m,t}\\ \end{bmatrix}$
+where $x_{i,j}$ is the value of feature $i$ at time $j$.
 Accompanying the asset matrix is the market data for the same $t$ time steps, represented as a matrix $Z\in\mathbb{R}^{g\times t}$, where $g$ is the number of market features, that is
 $Z=\begin{bmatrix}z_{1,1}&z_{1,2}&\cdots&z_{1,t}\\ z_{2,1}&z_{2,2}&\cdots&z_{2,t}\\ \vdots&\vdots&\ddots&\vdots\\ z_{g,1}&z_{g,2}&\cdots&z_{g,t}\\ \end{bmatrix}$
-Our goal is to predict the distribution of log-returns of the asset for the future time steps from $t+1$ to $t+n$, where $n$ represents both the number of time steps used for the predictions and the forecast horizon.
-The vector of log returns $\vec{r}\in\mathbb{R}^{n}$ is defined as:
-$\vec{r}=\begin{bmatrix}r_{t+1}\\ r_{t+2}\\ \vdots\\ r_{t+n}\\ \end{bmatrix}$
-where each $r_{t+k}$ represents the log return at time step $t+k$ for $k=1,\dots,n$.
-Given the input features $X$ and $Z$, the objective is to model and predict the distribution of future log returns
+where $z_{i,j}$ is value of market asset $i$ at time $j$.
+Our goal is to estimate the distribution of the log returns of the asset for the future time steps from $1$ to $t$, where $t$ represents both the number of time steps used for the estimation and the forecast horizon.
+The vector of log returns $\vec{r}\in\mathbb{R}^{t}$ is defined as:
+$\vec{r}=\begin{bmatrix}r_{1}\\ r_{2}\\ \vdots\\ r_{n}\\ \end{bmatrix}.$
+Given the input features $X$ and $Z$, the objective is to model and estimate the distribution of log returns, that is,
 $\vec{r}\sim f(X,Z)$
-where $f(X,Z)$ represents the predictive model for the log returns based on the asset and market features.
+where $f(X,Z)$ represents the model for the estimation of log returns based on the asset and market features.
 3.2 Model Architecture
-To model the distribution of future log returns, we aim to predict the quantiles $\vec{\tau}$ of the return distribution over the prediction horizon.
-Utilizing kernel density estimation these can later be converted into distributions.
+To model the distribution of log returns, our objective is to estimate the quantiles $\vec{\tau}$ of the log return distribution over the estimation horizon.
+Utilizing density estimation, these can later be converted into distributions.
 The quantile vector is defined as
-$\vec{\tau}=\begin{bmatrix}\tau_{1}\\ \tau_{2}\\ \vdots\\ \tau_{\mathcal{T}}\\ \end{bmatrix}$
-where $\mathcal{T}$ denotes the number of quantiles, the quantiles used can be found in F.
-For this task, we utilize Long Short-Term Memory (LSTM) networks, a type of recurrent neural network (RNN) [16] that are well-suited for sequential data modeling due to their ability to capture long-term dependencies and are less prone to issues such as vanishing gradients.
-Unlike many other time-series models, LSTMs can flexibly handle variable-length sequences, making them appropriate for our task as we want a variable amount of sequence lengths.
-Following the methodology outlined in [2], we adopt a two-stage approach to predict the distribution of future log returns.
-3.2.1 Stage 1: Asset-Specific Quantile Prediction
-In the first stage, the model focuses on predicting the quantiles of future normalized returns $\vec{\tilde{r}}$ by leveraging the asset-specific features $X$ and the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group.
-This is done by predicting the quantiles $Q_{\vec{\tilde{r}}}^{\tau}$ of the volatility-normalized returns $\vec{\tilde{r}}$, where
+$\vec{\tau}=\begin{bmatrix}\tau_{1}\\ \tau_{2}\\ \vdots\\ \tau_{\mathcal{T}}\\ \end{bmatrix}$ (1)
+where $\mathcal{T}$ denotes the number of quantiles.
+For the purpose of this research, we chose quantiles that represent broad ranging characteristics of a distribution, including quartiles and deciles as well as more fine grained intervals at the lower and upper end of the distribution to capture heavy tails.
+The quantiles used in this research can be found in A.
+For this task, we use long- and short-term memory (LSTM) networks, a type of recurrent neural network (RNN) [26], which are well suited for sequential data modeling due to their ability to capture long-term dependencies and are less prone to issues such as vanishing gradients.
+LSTMs maintain two states at each time step: the cell state $\vec{c}_{t}$ and the hidden state $\vec{h}_{t}$, which are updated using gating mechanisms.
+In addition there are the input and output gates, $\vec{i}_{t}$ and $\vec{o}_{t}$, respectively.
+Specifically, the updates are governed by the equations:
+$\vec{i}_{t}$ $=\sigma(W_{i}\vec{x}_{t}+U_{i}h_{t-1}+\vec{b}_{i}),$
+$\vec{f}_{t}$ $=\sigma(W_{f}\vec{x}_{t}+U_{f}\vec{h}_{t-1}+\vec{b}_{f}),$
+$\vec{o}_{t}$ $=\sigma(W_{o}\vec{x}_{t}+U_{o}\vec{h}_{t-1}+\vec{b}_{o}),$
+$\tilde{\vec{c}}_{t}$ $=\tanh(W_{c}\vec{x}_{t}+U_{c}\vec{h}_{t-1}+\vec{b}_{c}),$
+$\vec{c}_{t}$ $=\vec{f}_{t}\odot\vec{c}_{t-1}+\vec{i}_{t}\odot\tilde{\vec{c}}_{t},$
+$\vec{h}_{t}$ $=\vec{o}_{t}\odot\tanh(\vec{c}_{t}),$
+where $\sigma(\cdot)$ and $\tanh(\cdot)$ are the activation functions of the sigmoid and hyperbolic tangent, respectively, and $\odot$ denotes the multiplication of elements.
+Unlike many other time-series models, LSTMs can flexibly handle variable-length sequences, making them appropriate for our task as we are working with a variable number of sequence lengths.
+Loosely following the methodology outlined in [12], we adopt a two-stage approach to estimate the distribution of future log returns.
+3.2.1 Stage 1: Asset-Specific Quantile Estimation
+In the first stage, the model focuses on estimating the quantiles of future normalized returns $\vec{\tilde{r}}$ by leveraging the asset-specific features $X$.
+This is done by estimating the quantiles $Q_{\vec{\tilde{r}}}^{\tau}$ of the volatility-normalized returns $\vec{\tilde{r}}$, where
 $\vec{\tilde{r}}=\frac{\vec{r}}{\vec{\bar{\sigma}}}$
 with
-$\vec{r}=\begin{bmatrix}r_{t+1}\\ \vdots\\ r_{t+n}\end{bmatrix}\quad\textrm{and}\quad\vec{\bar{\sigma}}=\begin{bmatrix}\bar{\sigma}_{t+1}\\ \vdots\\ \bar{\sigma}_{t+n}\end{bmatrix}$
-Here, $\vec{\bar{\sigma}}$ represents the average volatility of assets in the same group at time step $t+k$, for $k=1,\dots,n$, where $n$ is the prediction horizon.
-The average volatility $\bar{\sigma}_{t}$ is calculated as the exponentially weighted moving average (EWMA) of the log returns across all $K$ assets in the group, with a decay factor of 0.94, in line with standard practices and the methodology from [2], formally:
+$\vec{r}=\begin{bmatrix}r_{1}\\ \vdots\\ r_{t}\end{bmatrix}\quad\textrm{and}\quad\vec{\bar{\sigma}}=\begin{bmatrix}\bar{\sigma}_{1}\\ \vdots\\ \bar{\sigma}_{t}\end{bmatrix}$
+Here, $\vec{\bar{\sigma}}$ represents the average volatility of assets in the same group at some time step, where $t$ is the estimation horizon.
+The average volatility $\bar{\sigma}_{t}$ is calculated as the exponentially weighted moving average (EWMA) of the log returns across all $K$ assets in the group, with a decay factor of 0.94, in line with standard practices and the methodology from [12].
+Formally the expression is
 $\sigma_{t}^{2}=\lambda\sigma_{t-1}^{2}+(1-\lambda)r_{t-1}^{2}$
-with $\lambda$ being the decay factor, in our case 0.94, and $\bar{\sigma}_{t}$ is calculated as:
+where $\lambda=0.94$ is the decay factor. Here, $\bar{\sigma}_{t}$ is calculated as
 $\bar{\sigma}_{t}=\frac{1}{K}\sum_{i=1}^{K}\sigma_{i,t}$
-where $\sigma_{i,t}$ is the volatility of asset $i$ at time $t$. We can formally represent this stage as:
-$f_{1}(X)=Q_{\vec{\tilde{r}}}^{\tau}$
+where $\sigma_{i,t}$ is the volatility of asset $i$ at time $t$. We represent this first stage with
+$f_{1}(X)=Q_{\vec{\tilde{r}}}^{\tau},$
+that is, a function of the input features gives the quantiles of the volatility-normalized log returns.
 3.2.2 Stage 2: Market Data Scaling
-In the second stage, the model incorporates the broader market data features $Z$ to refine the predictions from the first stage.
-The market data is used to create a scaling factor $\vec{\tilde{\sigma}}$, which adjusts the asset-specific quantile predictions based on market conditions.
-The scaling factor $\vec{\tilde{\sigma}}$ is generated by applying a transformation to the market feature matrix $Z$, capturing the influence of market conditions on future returns.
-Additionally, the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group is used to further scale the predicted quantiles.
-This results in the final quantile prediction for the asset’s log returns:
-$Q_{\vec{r}}^{\tau}=\vec{\hat{\sigma}}\cdot\vec{\bar{\sigma}}\cdot Q_{\vec{\tilde{r}}}^{\tau}$
+In the second stage, the model incorporates the broader market data features $Z$ to refine the estimation from the first stage.
+The market data is used to create a scaling factor $\vec{\tilde{\sigma}}$, which adjusts the asset-specific quantile estimation based on market conditions.
+The scaling factor $\vec{\hat{\sigma}}$ is generated by applying a transformation to the market feature matrix $Z$, capturing the influence of market conditions on future returns.
+This results in the final quantile estimation for the asset’s log returns:
+$Q_{\vec{r}}^{\tau}=\vec{\hat{\sigma}}\cdot Q_{\vec{\tilde{r}}}^{\tau}$ (2)
+where $Q_{\vec{r}}^{\tau}$ represents the final estimated quantile for the asset’s log returns, $\vec{\bar{\sigma}}$ is the scaling factor derived from the market data features $Z$, which adjusts the asset-specific estimations based on the broader market environment, and $Q_{\vec{\tilde{r}}}^{\tau}$ is the asset-specific quantile estimation from Stage 1, normalized by the asset’s volatility.
+We can formally represent this stage as
+$f_{2}(Z)=\vec{\hat{\sigma}},$
+that is, a function of the market data gives a scalar value to scale the quantiles from $f_{1}(X)$.
+Now let $f_{\text{DE}}$ be a Quantile-Smooth Density Estimation.
+This method takes quantiles and their associated probabilities to generate a smooth probability density function (PDF) while ensuring that the CDF is monotonic, bounded, and normalized.
+It smooths the density to prevent spikes, enforces non-negativity, and normalizes the PDF to integrate to one, providing a well-behaved representation of the underlying distribution; for more details see the appendix LABEL:app:DE.
+This leads to the final model which is given by the representation
+$\vec{r}\sim f(X,Z)=f_{\text{DE}}\left[f_{1}(X)\cdot f_{2}(Z)\right]$
+where $f_{\text{DE}}$ represents the Quantile-Smooth Density Estimation. We refer to this model as qLSTM.
+Hybrid model
+We also propose a hybrid model, which we refer to as qHybrid, where after training the qLSTM model, we compute the average of the quantile of the model and a Gaussian fit of the log returns in the lookback period, $-t$ to 00 used in $X$.
+Formally, we define this as
+$Q_{\mathcal{H}}^{\tau}=\frac{Q_{\mathcal{N}}^{\tau}+Q_{\vec{\tilde{r}}}^{\tau}}{2},$ (3)
+where $Q_{\mathcal{N}}^{\tau}$ represents the Gaussian quantiles across the $\vec{\tau}$ defined in Eq. (1), computed as
+$Q_{\mathcal{N}}^{\tau}=[\mu+\sigma\Phi^{-1}(\tau_{i})]_{i=1}^{\mathcal{T}},$
+with $\Phi(x)$ being the standard normal cumulative distribution function:
 […]
-$f_{2}(Z)=\vec{\hat{\sigma}}$
-Giving the final model the representation:
-$\vec{r}\sim f(X,Z)=f_{\text{KDE}}(f_{1}(X)\cdot\vec{\bar{\sigma}}\cdot f_{2}(Z)$
-where $f_{\text{KDE}}$ represents the kernel density estimations, for more details see Appendix D.
+Here, $\mu$ and $\sigma$ denote the mean and standard deviation of the observed returns in the lookback period, from $-t$ to 00.
 3.2.3 Loss Function
 Following the methodology of [12], we use a custom loss function based on quantile regression to train the model.
-The original loss function is designed to minimize the error between the true returns and the predicted quantiles.
+The original loss function is designed to minimize the error between the true returns and the estimated quantiles.
 It is defined as
-$L_{\mathcal{T}}=\frac{1}{BK}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i}-\hat{Q}_{r_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i}-\hat{Q}_{\tilde{r}_{i}}(\tau)\right)\right]$ (1)
-where $\mathcal{T}$ is the set of quantiles used for the prediction, $B$ is the batch size (i.e., the number of assets in each batch), $K$ is the number of quantiles in the set $\mathcal{T}$, $r_{i}$ is the raw return of the asset, $\tilde{r}_{i}$ is the standardized (volatility-normalized) return of the asset and $\hat{Q}_{r_{i}}(\tau)$ and $\hat{Q}_{\tilde{r}_{i}}(\tau)$ are the predicted quantiles of the raw and standardized returns for quantile $\tau$, respectively Furhtermore, $\rho_{\tau}(\xi)$ is the quantile loss function
+$\mathcal{L}_{\tau}=\frac{1}{BK}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i}-\hat{Q}_{r_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i}-\hat{Q}_{\tilde{r}_{i}}(\tau)\right)\right]$ (4)
+where $\mathcal{T}$ is the set of quantiles used for the estimation, $B$ is the batch size (i.e., the number of assets in each batch), $K$ is the number of quantiles in the set $\mathcal{T}$, $r_{i}$ is the raw return of the asset, $\tilde{r}_{i}$ is the standardized (volatility-normalized) return of the asset and $\hat{Q}_{r_{i}}(\tau)$ and $\hat{Q}_{\tilde{r}_{i}}(\tau)$ are the estimated quantiles of the raw and standardized returns for quantile $\tau$, respectively Furhtermore, $\rho_{\tau}(\xi)$ is the quantile loss function
 $\rho_{\tau}(\xi)=\begin{cases}\tau\cdot\xi,&\text{if }\xi\geq 0\\ (\tau-1)\cdot\xi,&\text{if }\xi<0\end{cases}$
-where $\xi$ represents the error (residual) between the true return value and the predicted quantile.
+where $\xi$ represents the error (residual) between the true return value and the estimated quantile.
 3.2.4 Modified Loss Function
-Our proposed model generalizes the model in [2] by generating multiple predictions over a look-ahead window, and we therefore modify the loss function in 1 to account for each future time step $t\in\{1,2,\dots,T\}$ within the prediction horizon.
+Our proposed model generalizes the model in [12] by generating a singe estimation for log returns over a look-ahead window, and we therefore modify the loss function in Eq.(4) to account for each future time step $t\in\{1,2,\dots,T\}$ within the estimation horizon.
 The modified loss function is defined as
-$L_{\mathcal{T}}=\frac{1}{BTK}\sum_{\tau\in\mathcal{T}}\sum_{t=1}^{T}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i,t}-\hat{Q}_{\vec{r}_{i,t}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i,t}-\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)\right)\right]$
-where $\mathcal{T}$ is the number of future time steps (i.e., the prediction horizon) in the look-ahead window, $r_{i,t}$ and $\tilde{r}_{i,t}$ are the raw and standardized returns of the asset at future time step $t$, and $\hat{Q}_{\vec{r}_{i,t}}(\tau)$ and $\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)$ are the predicted quantiles of the raw and standardized returns for asset $i$ at future time step $t$.
-This modification allows us to compute the loss over the entire prediction window, training the model to minimize the quantile loss for each future time step within the horizon.
+$\mathcal{L}_{\mathcal{T}}=\frac{1}{BTK}\sum_{\tau\in\mathcal{T}}\sum_{t=1}^{T}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i,t}-\hat{Q}_{\vec{r}_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i,t}-\hat{Q}_{\vec{\tilde{r}}_{i}}(\tau)\right)\right]$
+where $T$ is the number of future time steps (i.e., the estimation horizon) in the look-ahead window, $r_{i,t}$ and $\tilde{r}_{i,t}$ are the raw and standardized returns of the asset at future time step $t$, and $\hat{Q}_{\vec{r}_{i,t}}(\tau)$ and $\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)$ are the estimated quantiles of the raw and standardized returns for asset $i$ at future time step $t$.
+This modification allows us to compute the loss over the entire estimation window, training the model to minimize the quantile loss for each future time step within the horizon.
 The inclusion of both raw and standardized returns helps ensure that the model learns both absolute and volatility-adjusted relationships.
@@
 Within stocks there are three sub-asset classes for the index the stocks are listed on, S&P 500, Euro Stoxx 50, and Nikkei 225.
-Full details on the selected assets as well as summary statistics, can be found in Appendix A.
+Full details on the selected assets as well as summary statistics, can be found in Appendix C.1.
 Table 1: Summary of selected assets
@@
 3.3.2 Feature Engineering
-Additional features are computed for each asset to enhance the model’s ability to capture various market dynamics.
+Additional features are computed for each asset to enhance the models’ ability to capture various market dynamics.
 Table 2 shows an overview of the engineered features.
@@
 This data includes stock indices, commodities, other indices, and treasury yields.
-For more details on market data, see Appendix B.
+For more details on market data, see Appendix C.2.
 Table 3: Overview of market data sources
@@
 Treasury Yields 10-year and 3-month Treasury yields.
-3.3.4 Data Split
-The data is split into training, validation, and test sets based on the time periods as follows: Training set: All available data up to the end of 2017.
-Validation set: Data from 2018 to the end of 2019.
-Test set: Data from 2019 to 2024, which includes the COVID-19 pandemic and the 2022 bear market.
-To create varying sequence lengths, each asset’s data is randomly divided into sequences with lengths sampled uniformly from the range $[1,30]$ days.
 3.4 Synthetic Data
@@
 Additionally, we generate $g$ synthetic datasets for the market data, where $g$ represents the number of market features.
-One feature is designed to have a correlation of approximately 0.7 with the synthetic returns to simulate market co-movement For more details see Appendix E.
+One feature is designed to have a correlation of approximately 0.7 with the synthetic returns to simulate market co-movement For more details see Appendix C.3.
 3.4.1 Hyperparameter Optimization
-The model’s hyperparameters are optimized using Bayesian optimization, with separate sets of hyperparameters for the asset-specific model (Stage 1) and the market data model (Stage 2).
+The models’ hyperparameters are optimized using Bayesian optimization, with separate sets of hyperparameters for the asset-specific model (Stage 1) and the market data model (Stage 2) [31].
 Each set of hyperparameters is sampled independently.
 The common hyperparameters and the specific hyperparameters for each stage are detailed in Table 4.
-Table 4: Hyperparameters for Asset-Specific and Market Data Models
+The optimal hyperparameters found for qLSTM and subsequently qHybrid -because they use the same hyperparameters- can be found in Appendix D.1.
+Table 4: Hyperparameters for qLSTM
 Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
@@
 L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
-Use layer normalization Boolean: $\{1,0\}$
+Use layer
+normalization Boolean: $\{1,0\}$
 LSTM layers $l_{\text{LSTM, asset}}\in[1,5]$ $l_{\text{LSTM, market}}\in[1,5]$
@@
 3.5 Experimental Setup
-Our proposed model -which will be referred to as qLSTM- is evaluated and compared against the original model architecture [2] -which will be called qDense-.
-It should be noted that the comparison is not entirely fair as the original model was trained under different circumstances and utilized different data.
-As the qDense model has a fixed forecast window of 22 days, the qLSTM model will only utilize a 22-day-ahead window as well.
-This means that on the test data, the assets are iterated sequentially taking 22 entries at a time, and then the loss of the 22-day-quantiles of both models is compared using the loss function described in the next section.
-To act as a baseline for both models, we compare them to a simple linear quantile regression (LQR) model  [^1] The LQR model is trained and tested using the same data as qDense [22].
+Our proposed models qLSTM, defined in Eq.
+(2) and qHybrid, found in Eq.
+(3) is evaluated and compared against the model architecture from [12] -which will be called qDense.
+It should be noted that the comparison is not entirely fair as the qDense model was trained under different circumstances and utilized different data.
+However, to ensure fair comparison as much as we can, in this paper, we always train the models on the same setup.
+As the qDense model has a fixed forecast window of 22 days, the qLSTM and qHybrid models will only utilize a 22-day-ahead window as well.
+This means that on the test data, the assets are iterated sequentially taking 22 entries at a time, and then the loss of the 22-day quantiles of both models is compared using the loss function described in the next section.
+To act as a baseline for all models, we compare them to a simple linear quantile regression (LQR) model  [^1] The LQR model is trained and tested using the same data as qDense [32].
+3.5.1 Data Split
+The data is split into training, validation, and test sets based on the time periods as follows
+  • 1.
+    Training set: All available data up to the end of 2017.
+  • 2.
+    Validation set: Data from 2018 to the end of 2019.
+  • 3.
+    Test set: Data from 2019 to 2024, which includes the COVID-19 pandemic and the 2022 bear market.
+To create varying sequence lengths, each asset’s data is randomly divided into sequences with lengths sampled uniformly from the range $[15,30]$ days.
 3.5.2 Training Procedure
-The model is trained using the Adam optimizer with a fixed number of 100 epochs.
+The qLSTM model is trained using the Adam optimizer with a fixed number of 100 epochs.
 Early stopping is employed with a patience parameter of 10, meaning that if the validation loss does not improve for 10 consecutive epochs, training is halted.
 The weights that generated the lowest loss on the validation data are chosen.
-3.6 Evaluation Metric
-The primary evaluation metric is the quantile loss function, which has been slightly modified to disregard the volatility normalization as well as using only a single return for the evaluation.
+The training details for qDense can be found in appendix D.2.
+3.6 Evaluation Metrics
+To evaluate the performance of our propsoed modesl, qLSTM and qHybrid, we use several measures, which assess the distribution accuracy in different ways as described below.
+3.6.1 Quantile Loss
+The primary evaluation metric for comparing qLSTM, qHybrid, with qDense and LQR is the quantile loss function, which has been slightly modified to ignore volatility normalization and to use only one return for the evaluation.
 The quantile loss for a given asset is defined as
 $L_{\mathcal{T}}=\frac{1}{KN}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{T}\rho_{\tau}\left(r_{t}-\hat{Q}_{r_{t}}(\tau)\right),$
-where $\mathcal{T}$ is the set of quantiles used for prediction, and $K=|\mathcal{T}|$ is the number of quantiles, $T$ is the number of time steps in the test data for the given asset, $r_{t}$ is the raw return of the asset at time step $t$, and $\hat{Q}_{r_{t}}(\tau)$ is the predicted quantile of the raw return of the asset for quantile level $\tau$ at time step $t$.
+where $\mathcal{T}$ is the set of quantiles used for estimation, and $K=|\mathcal{T}|$ is the number of quantiles, $T$ is the number of time steps in the test data for the given asset, $r_{t}$ is the raw return of the asset at time step $t$, and $\hat{Q}_{r_{t}}(\tau)$ is the estimated quantile of the raw return of the asset for quantile level $\tau$ at time step $t$.
 $\rho_{\tau}(\xi)$ is the quantile loss function, defined as:
 $\rho_{\tau}(\xi)=\begin{cases}\tau\cdot\xi,&\text{if }\xi\geq 0\\ (\tau-1)\cdot\xi,&\text{if }\xi<0\end{cases}$
-with $\xi=r_{i,t,j}-\hat{Q}_{r_{i,t,j}}(\tau)$ representing the residual (error) between the true return and the predicted quantile.
+with $\xi=r_{i,t,j}-\hat{Q}_{r_{i,t,j}}(\tau)$ representing the residual (error) between the true return and the estimated quantile.
+3.6.2 Distribution Metrics
+For the following metrics, we utilize the fact that we estimate the quantiles across multiple observations. We used a forecast window of 30 to get the best estimate for the empirical distribution.
+RMSE for Moments
+We further evaluate the qLSTM and qHybrid on root mean squared error (RMSE) in moments estimation and compare them to the assumption that the moments are unchanged.
+, We evalute them with respect to the first four moments, mean ($\mu$), standard deviation ($\sigma$), kurtosis ($\gamma$), and skew ($\kappa$).
+The error is computes using the formula
+$\text{RMSE}_{j}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\lambda_{i,j}-\hat{\lambda}_{i,j})^{2}}$
+where $N$ is the test sample, $\lambda$ is the realized moment, $\hat{\lambda}$ is the estimated moment and $j$ is the moment number.
+Wasserstain Distance
+To evaluate the alignment between the estimated distribution of qLSTM, qHybrid, and the observed returns, we incorporate the Wasserstein distance as an additional metric and compare it against the normal distribution as a benchmark.
+The Wasserstein distance measures the cost of transforming the estimated distribution into the realized distribution.
+For a estimated(PDF $p(x)$, discretized over grid points $\{x_{i}\}_{i=1}^{N}$, and realized returns $\{r_{i}\}_{i=1}^{M}$, the Wasserstein distance is defined as
+[…]
+where $P$ is the estimated empirical distribution and $Q$ is the empirical distribution of realized returns.
+In practice, we approximate this by sampling $M$ estimated values from $p(x)$ and matching them to the $M$ realized returns, resulting in
+$W_{1}(P,Q)=\frac{1}{M}\sum_{i=1}^{M}\left|x_{i}^{*}-r_{i}\right|,$
+where $\{x_{i}^{*}\}_{i=1}^{M}$ are the sampled values from the estimated PDF, and $\{r_{i}\}_{i=1}^{M}$ are the sorted realized returns.
+Lower Wasserstein distances indicate better alignment between the estimated and realized distributions.
+Continuous Ranked Probability Score
+The Continuous Ranked Probability Score (CRPS) evaluates the alignment between the estimated CDF and the observed values.
+It measures the integrated squared difference between the estimated CDF and the empirical CDF of the observations.
+The CRPS for a single observation is defined as
+[…]
+where $F(x)$ is the estimated CDF for a given observation, $x_{\text{obs}}$ is the observed value, and $\mathbf{1}(x\geq x_{\text{obs}})$ is the indicator function that equals 1 if $x\geq x_{\text{obs}}$ and 0 otherwise.
+For discrete predictions, the CRPS can be approximated as
+$\text{CRPS}(F,x_{\text{obs}})=\frac{1}{N}\sum_{i=1}^{N}\left(F(x_{i})-\mathbf{1}(x_{i}\geq x_{\text{obs}})\right)^{2},$
+where $\{x_{i}\}_{i=1}^{N}$ represents the discretized grid over which $F(x)$ is evaluated. The mean CRPS across all samples in the test set is calculated as:
+Mean $\text{Mean CRPS}=\frac{1}{M}\sum_{j=1}^{M}\text{CRPS}(F_{j},x_{\text{obs},j}),$
+where $M$ is the total number of observations, $F_{j}$ is the estimated CDF for observation $j$, and $x_{\text{obs},j}$ is the corresponding observed value.
+This metric provides a comprehensive evaluation of the probabilistic calibration of the model by comparing the estimated distribution with the observed results.
+Lower CRPS values indicate better alignment between the estimated and observed distributions.
+Deviation from Quantile Level (VaR Distance)
+To evaluate the calibration of the qLSTM and qHybrid models to estimate tail risks, we calculate the deviation from the desired quantile level at specific quantiles $\tau$ and compare it against a Gaussian baseline.
+This metric measures how far the observed violation rate deviates from the expected quantile level, with smaller deviations indicating better calibration.
+The Violation Rate (VR) is defined as
+$\text{VR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(r_{i}<\hat{Q}_{r}(\tau)),$
+where $N$ is the total number of observations, $r_{i}$ is the observed return at time $i$,$\hat{Q}_{r}(\tau)$ is the edstimated quantile (VaR) at level $\tau$, and $\mathbf{1}(x<y)$ is the indicator function that equals 1 if $x<y$, and 0 otherwise.
+The deviation from the quantile level $\tau$ is then calculated as
+$\text{Deviation}=\left|\text{VR}-\tau\right|,$
+where VR is the Violation Rate, and $\tau$ is the expected quantile level (e.g., 0.05, 0.01, 0.00075). This deviation is computed for both the estimated quantiles and a Gaussian baseline.
+The resulting metric provides a measure of the model’s calibration quality.
+A perfectly calibrated model would have $\text{VR}=\tau$, leading to a deviation of 0.
+Larger deviations indicate underestimation or overestimation of tail risk.
 4 Results & Discussion
-4.1 Hyper parameter Optimization
-Table 5 shows the optimal hyperparameters found in the 100 trials on market data, the final model has a total of $228.390$ trainable parameters.
-The optimal hyperparameters of the qDense model can be found in Appendix C.
-Table 5: Final Hyperparameters for the qLSTM model, rounded down to the 4th decimal place
-Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
-Batch size 256
-Learning rate 0.0004
-Normalization window 34
-Raw LSTM layers 4 1
-LSTM units 64 128
-Dense layers 1 1
-Dense units per layer 128 64
-Dropout rate 0.1023
-Market activation None
-Hidden activation Leaky ReLU
-Use layer normalization Yes (1)
-L1 regularization 0.0009
-L2 regularization 0.0010
-4.2 Model Performance Evaluation
-Table 6: Comparison of qDense, qLSTM, and LQR across various asset types. Percentage differences are calculated using mean percentage difference.
-LQR qDense qLSTM qDense - qLSTM LQR - qLSTM
-Commodities 1.4850 0.6226 0.3319 0.2908 (60.93%) 1.1532 (126.94%)
-Crypto-currencies 2.5176 1.3605 0.9292 0.4313 (37.68%) 1.5884 (92.17%)
-S&P 500 1.4832 0.5969 0.3284 0.2684 (58.02%) 1.1547 (127.48%)
-Nikkei 225 1.4805 0.6187 0.3298 0.2889 (60.92%) 1.1507 (127.13%)
-Euro Stoxx 50 1.4803 0.5983 0.3326 0.2656 (57.07%) 1.1477 (126.61%)
-Currency Pairs 1.2086 0.1597 0.1156 0.0441 (32.05%) 1.0930 (165.09%)
-Normal Synthetic 0.2400 0.7003 0.3549 0.3454 (65.46%) -0.1149 (-38.62%)
-Log Normal Synthetic 0.4163 0.7075 0.3713 0.3362 (62.33%) 0.0450 (11.43%)
-Gamma Synthetic 0.5124 0.7392 0.3805 0.3586 (64.06%) 0.1318 (29.53%)
-Uniform Synthetic 0.2188 0.5683 0.2993 0.2691 (62.03%) -0.0805 (-31.06%)
-Total Market Mean 1.6092 0.6594 0.3946 0.2649 (50.26%) 1.2146 (121.23%)
-Total Synthetic Mean 0.3469 0.6788 0.3515 0.3273 (63.54%) -0.0046 (-1.32%)
-Total Mean 1.1043 0.6672 0.3773 0.2898 (55.50%) 0.7269 (98.13%)
-The results of the model evaluation across various assets and synthetic datasets, measured in quantile loss, are shown in Table LABEL:tab:final-results.
-Both the qLSTM and qDense models outperform the Linear Quantile Regression (LQR) model on market data.
-However, LQR manages to outperform both models on some of the synthetic data, as illustrated in Table LABEL:tab:final-results.
-The qLSTM model consistently outperforms the qDense model, with an average performance improvement of 50% on real-world datasets, 64% on synthetic datasets, and 55% overall.
-Figure 1: The PDFs generated by qLSTM and qDense the first 100 out-of-sample observations of BTC-USD
-On real-world data, the qLSTM model demonstrates a significant performance boost.
-For instance, when applied to S&P 500 stocks, the qLSTM model achieves a quantile loss of 0.3284, compared to 0.5969 for the qDense model and 1.4805 for LQR—representing a 58% and 127% improvement, respectively.
-Similarly, for cryptocurrencies, the qLSTM model reduces quantile loss by 37% compared to qDense and 92% compared to LQR, showcasing its robustness in handling volatile and unpredictable assets.
-Figure 1 shows the distribution of quantile losses for qLSTM and qDense over time for the real-world data.
-The qDense distribution undergoes greater shape variation and appears more symmetrical, while the qLSTM model captures more complexity with less variation in shape, indicating its ability to model intricate distribution patterns.
-This suggests that qLSTM captures dynamic distributions more effectively, whereas qDense tends to produce distributions closer to normality.
-This is further illustrated in Figure 2, where a single slice of the distributions shows qDense’s smoother, nearly normal distribution, while qLSTM displays a more intricate and asymmetrical shape, which is better aligned with real data.
-Figure 2: The PDFs generated by qLSTM and qDense on a randomly chosen out-of-sample observation of BTC-USD
-When evaluated on synthetic data, the qLSTM model shows greater improvement over qDense, but still falls short of LQR on average.
-qLSTM outperforms LQR on log-normal and gamma distributions, improving quantile loss by 11% and 30%, respectively, demonstrating its ability to capture asymmetrical distributions effectively.
-However, on the normal and uniform distributions, which are largely symmetrical, LQR outperforms qLSTM by 39% and 31%, respectively.
-This is somewhat expected, given the nice, symmetric behavior of these distributions.
-Figure 3 shows the first 100 samples of synthetically generated normally distributed data, where neither distribution changes significantly in shape, with LQR showing no notable variation, as expected for i.i.d.
-data.
-This reflects how well the models capture the structure of synthetic data.
-Although LQR performs better on symmetrical distributions, qLSTM still generates somewhat asymmetrical distributions that approach normality.
-Figure 4 presents a single slice of the distributions produced by LQR and qLSTM, highlighting qLSTM’s simpler distribution in comparison to real-world data, though still exhibiting slight asymmetry.
-Interestingly, LQR shows a more complex shape in this case, despite achieving lower quantile loss than qLSTM on this dataset.
-Figure 3: The PDFs generated by qLSTM and LQR on the first 100 synthetic normally distributed data
-Figure 4: The PDFs generated by qLSTM and LQR on a randomly chosen observation of synthetic normally distributed data
+In this section we present the performance of our proposed models and compare them to baseline models using various metrics.
+4.1 Quantile Loss Performance
+Table 5: Quantiles loss for synthetic and non-synthetic data across models.
+Quantile loss
+Data Type Asset Class LQM qDense qLSTM qHybrid
+Real World Commodities 1.4815 0.4136 0.3223 0.3693
+Cryptocurrencies 2.4931 2.2843 1.0222 1.2323
+S&P 500 1.4811 0.4255 0.3150 0.3572
+Nikkei 225 1.4739 0.3991 0.3141 0.3552
+Euro Stoxx 50 1.4772 0.4184 0.3124 0.3520
+Currency Pairs 1.2003 0.0788 0.1528 0.0973
+Total Real World 1.6012 0.6699 0.4065 0.4606
+Synthetic Normal 0.5029 0.9338 0.4201 0.4572
+Log Normal 0.2115 0.7229 0.3230 0.1565
+Gamma 0.3260 0.8268 0.3642 0.2781
+Uniform 0.3366 0.6864 0.2940 0.3030
+Total Synthetic 0.3443 0.7925 0.3503 0.2987
+Total 1.0984 0.7190 0.3840 0.3958
+The results in Table 5 highlight the effectiveness of different models (LQM, qDense, qLSTM, and qHybrid) across real-world and synthetic datasets.
+Overall, the qLSTM model demonstrates superior performance, achieving the lowest quantile loss in most real-world asset classes, including commodities, cryptocurrencies, and equity indices such as the S&P 500, Nikkei 225, and Euro Stoxx 50.
+Notably, for cryptocurrencies, LSTM reduces the quantile loss to 1.0222 compared to 2.2843 for the Dense model and 2.4931 for the Linear model, showcasing its ability to effectively capture the volatility and non-linearity inherent in this asset class.
+However, for currency pairs, the Dense model performs best, achieving a quantile loss of 0.0788, which may indicate that simpler architectures can suffice for less complex financial data.
+For synthetic data, the qHybrid model demonstrates an edge, particularly for skewed and heavy-tailed distributions such as Log Normal and Gamma.
+It achieves the lowest quantile losses of 0.1565 and 0.2781, respectively, suggesting that the combination of Gaussian assumptions with learned quantiles is advantageous in these scenarios.
+However, for simpler distributions such as Normal and Uniform, LSTM outperforms the qHybrid model, indicating its ability to generalize across different statistical patterns.
+In terms of overall performance, LSTM achieves the lowest total quantile loss for real-world data (0.4065), while the qHybrid model marginally outperforms it for synthetic datasets (0.2987 vs.
+0.3503).
+When considering all datasets, the LSTM model maintains the lowest overall total quantile loss (0.3840), reflecting its versatility and robustness.
+These findings underscore the strengths of the qLSTM models in handling real-world sequential financial data, particularly for volatile asset classes like cryptocurrencies and equity indices.
+The qHybrid model’s success with synthetic data highlights its potential for capturing diverse statistical properties, which could complement LSTM’s strengths in real-world applications.
+Overall, the two proposed models are able to learn the quantile distribution of sequential data.
+4.2 Distribution Metric Performance
+In this section, we explore the results of the distribution metrics for qLSTM and qHybrid against a baseline.
+Note for the sake of conciseness and relevance we omit the comparison against synthetic data and focus only on real world data.
+Table 6: Root mean square error of moments estimation compared against the assumption that moments are constant.
+Asset class
+Stat-istic Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
+$\mu$ qLSTM 0.0118 0.0010 0.0034 0.0041 0.0032 0.0034
+qHybrid 0.0119 0.0020 0.0035 0.0043 0.0033 0.0035
+Constant 0.0166 0.0013 0.0050 0.0059 0.0048 0.0050
+$\sigma$ qLSTM 0.0424 0.0287 0.0085 0.0091 0.0096 0.0068
+qHybrid 0.0400 0.0422 0.0088 0.0109 0.0112 0.0084
+Constant 0.0330 0.0021 0.0077 0.0101 0.0099 0.0074
+$\gamma$ qLSTM 1.1129 0.7186 0.7667 0.7770 0.8350 0.7422
+qHybrid 1.1133 0.7030 0.7496 0.7585 0.8145 0.7339
+Constant 1.6036 0.9410 1.0207 1.0598 1.1718 1.0596
+$\kappa$ qLSTM 8.1838 9.2768 8.9404 8.7593 8.7733 8.8889
+qHybrid 5.3220 4.4449 4.2748 4.0337 4.1521 4.1221
+Constant 4.3678 2.7381 2.8323 2.6574 3.1615 2.7176
+Table 6 presents the root mean square error (RMSE) for moment estimation across the models qLSTM, qHybrid, and Constant -the assumption that moments are constant throughout $-t$ to 00 and $1$ to $t$- and asset classes.
+For the mean $\mu$, the qLSTM model consistently achieves the lowest RMSE in all asset classes, highlighting its superior ability to accurately estimate the central tendency of the returns.
+While the qHybrid model performs comparably, it slightly trails the qLSTM in most categories.
+The Constant model demonstrates the highest RMSE, suggesting it struggles to capture the mean effectively, particularly for cryptocurrencies and equity indices.
+For the standard deviation $\sigma$, the performance varies.
+The qLSTM model outperforms others in most equity indices, commodities, and cryptocurrencies, reflecting its ability to capture volatility dynamics effectively.
+However, the Constant model achieves the lowest RMSE for currency pairs, suggesting that simpler methods may suffice for less complex financial assets.
+The qHybrid model provides competitive results but generally underperforms compared to qLSTM.
+The results of skewness $\gamma$ indicate that the qLSTM and qHybrid models perform closely, with the qHybrid model having a slight edge in most asset classes.
+This suggests that qHybrid models may have an advantage in capturing asymmetry in return distributions.
+However, the Constant model has significantly higher RMSE values, indicating poor performance in estimating skewness, especially for cryptocurrencies and equity indices.
+Finally, for kurtosis $\kappa$, the Constant model achieves the lowest RMSE values, suggesting it captures tail behavior more effectively in this metric.
+Both qLSTM and qHybrid models show higher RMSE values for kurtosis, which may indicate that their emphasis on other moments slightly compromises their accuracy in tail risk estimation.
+Nonetheless, qLSTM still maintains competitive performance, particularly for more volatile assets like cryptocurrencies.
+Table 7: Wasserstein distances for different models and asset classes. Lowest values are highlighted in bold.
+Asset class
+Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
+qLSTM 0.0260 0.0179 0.0079 0.0077 0.0081 0.0075
+qHybrid 0.0213 0.0194 0.0072 0.0079 0.0080 0.0071
+Gaussian 0.0221 0.0017 0.0074 0.0081 0.0079 0.0074
+Asset class
+Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
+qLSTM 0.4320 0.4477 0.4422 0.4424 0.4422 0.4421
+qHybrid 0.4179 0.4464 0.4331 0.4340 0.4338 0.4334
+Gaussian 0.3544 0.4793 0.4308 0.4366 0.4369 0.4349
+Table 8: CRPS values for different models, with the lowest values highlighted.
+Table 7 presents the Wasserstein distances for different models and asset classes, highlighting the alignment between the estimated and realized distributions.
+The qHybrid model achieves the lowest distances for most asset classes, particularly for cryptocurrencies (0.0213) and commodities (0.0072), showcasing its ability to provide a close match to the empirical distributions.
+However, for currency pairs (FX), the Gaussian model outperforms others with a distance of 0.0017, likely due to the simpler statistical patterns in this asset class.
+The qLSTM model remains competitive, delivering consistently low Wasserstein distances across asset classes, indicating its robustness in capturing the underlying return distributions.
+Table 8 evaluates the Continuous Ranked Probability Score (CRPS) for the models, which evaluates the probabilistic calibration of the estimated distributions.
+The qHybrid model performs best for most asset classes, including FX (0.4464), Euro Stoxx 50 (0.4340), S&P 500 (0.4338), and Nikkei 225 (0.4334), demonstrating its ability to provide well-calibrated probabilistic forecasts.
+For cryptocurrencies, the Gaussian model achieves the lowest CRPS (0.3544), suggesting its suitability for this highly volatile asset class.
+The qLSTM model offers competitive results, with scores close to the best-performing models, reflecting its strong calibration across diverse asset types.
+These results highlight the complementary strengths of the qHybrid and qLSTM models, with qHybrid excelling in calibration and qLSTM providing robust distributional alignment.
+Asset class
+Level Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
+0.05 qLSTM 0.1967 0.0499 0.0659 0.0539 0.0609 0.0549
+qHybrid 0.0777 0.0492 0.0481 0.0485 0.0506 0.0460
+Gaussian 0.0505 0.0505 0.0466 0.0537 0.0536 0.0469
+0.01 qLSTM 0.1085 0.0100 0.0206 0.0190 0.0208 0.0176
+qHybrid 0.0478 0.0100 0.0203 0.0220 0.0215 0.0198
+Gaussian 0.0298 0.0287 0.0262 0.0326 0.0322 0.0268
+0.00075 qLSTM 0.0353 0.0008 0.0028 0.0031 0.0024 0.0018
+qHybrid 0.0176 0.0008 0.0036 0.0056 0.0047 0.0029
+Gaussian 0.0139 0.0113 0.0104 0.0158 0.0157 0.0099
+Table 9: VaR values at different levels for various models, with the lowest values highlighted.
+Table 9 presents the Value at Risk (VaR) results for three models (qLSTM, qHybrid, and Gaussian) across asset classes at different confidence levels (0.05, 0.01, and 0.00075).
+At the 0.05 level, the Gaussian model achieves the lowest VaR for most asset classes, particularly for cryptocurrencies (0.0505) and commodities (0.0466).
+However, the qHybrid model demonstrates competitive performance, achieving the lowest VaR for FX (0.0492), Euro Stoxx 50 (0.0485), S&P 500 (0.0506), and Nikkei 225 (0.0460), indicating its strength in tail risk modeling across diverse assets.
+The qLSTM model performs reasonably well but tends to have higher VaR values, particularly for volatile assets like cryptocurrencies.
+At the 0.01 level, the qHybrid model shows robust performance across most asset classes, achieving the lowest VaR for commodities (0.0203) and S&P 500 (0.0215).
+The qLSTM model outperforms the other models for FX (0.0100), Euro Stoxx 50 (0.0190), and Nikkei 225 (0.0176), demonstrating its ability to capture extreme tail risks in these scenarios effectively.
+The Gaussian model remains competitive, particularly for cryptocurrencies, achieving a VaR of 0.0298, highlighting its utility in assets with higher volatility.
+At the extreme tail (0.00075 level), the qLSTM model consistently achieves the lowest VaR across most asset classes, including FX (0.0008), commodities (0.0028), Euro Stoxx 50 (0.0031), S&P 500 (0.0024), and Nikkei 225 (0.0018).
+This suggests that the qLSTM is particularly adept at capturing risks at the far ends of the distribution.
+The qHybrid and Gaussian models are slightly less effective at this level, though the Gaussian model shows competitive performance for cryptocurrencies (0.0139).
+These results emphasize the qLSTM’s strength in extreme tail risk estimation and the qHybrid model’s reliability across broader confidence levels.
 4.3 Limitations and Challenges
@@
 This issue arises primarily due to the lack of access to the original dataset used in their study and the absence of publicly available code for their implementation.
-Consequently, we were unable to replicate their exact methodology, which inherently limits the comparison between qLSTM approach and their dense model.
-Furthermore, the two approaches differ fundamentally in how they process data.
-The qLSTM model is specifically designed to handle sequence data, allowing it to capture patterns across variable-length sequences and predict multiple quantiles for future time steps simultaneously.
-In contrast, the dense model used in [2] operates on fixed-length windows and focuses on single-step quantile predictions.
-This introduces a degree of inconsistency in the comparison, as the qLSTM’s ability to process sequential information may provide it with a natural advantage in capturing long-term dependencies in asset returns.
-Another challenge emerged when comparing model performance on synthetic data.
-Due to differences in code constraints and hyperparameter configurations, the two models were not exposed to the same number of observations.
-Specifically, the qLSTM model used a normalization window of 34 days, while the qDense model operated on a longer window of 123 days.
-This discrepancy means that the qLSTM model was trained on a larger number of observations, increasing its potential for failure.
-However, despite being exposed to more opportunities for failure, the qLSTM model still outperformed the qDense model.
+Consequently, we were unable to replicate their exact methodology, which inherently limits the comparison between the qLSTM approach and their dense model.
+In terms of model performance, while the qLSTM model demonstrated superior results in tail risk estimation and overall robustness, its slightly higher Value at Risk (VaR) at broader confidence levels (e.g., 0.05) compared to the Gaussian and qHybrid models suggests room for improvement in capturing broader distributional patterns.
+Additionally, the qHybrid model, while effective in several scenarios, exhibited slightly higher RMSE for moments like kurtosis, indicating potential difficulties in accurately modeling extreme tails for certain asset classes.
+Finally, our results reveal some variability in model performance across asset classes.
+For example, while qLSTM excelled in cryptocurrencies and equity indices, the qDense and qHybrid models outperformed it in currency pairs and synthetic datasets with skewed distributions.
+These findings suggest that no single model is universally optimal, underscoring the importance of tailoring model architectures to specific data characteristics and use cases.
 5 Conclusion
-In this paper, we presented a novel model for predicting the distribution of asset returns using a quantile-based approach, extending prior research.
-By leveraging Long Short-Term Memory (LSTM) networks, we modeled the distribution of returns more dynamically and generalized these predictions across multiple asset classes, including synthetic data.
-Our approach diverged from previous models by incorporating asset-neutral features and utilizing market data to further refine the predictions.
-The results demonstrated significant improvements over a baseline linear quantile regression model, outperforming it by 98% across all datasets.
-Additionally, the LSTM approach outperformed a dense network model by over 50% on both real-world and synthetic data.
-This performance boost underscores the strength of our approach in capturing the complexities of financial return distributions, particularly for high-volatility assets like cryptocurrencies.
-Furthermore, the model’s ability to maintain strong performance across diverse synthetic datasets highlights its generalizability and flexibility.
-For future work, expanding the dataset to include more assets and asset classes—such as bonds, options, or credit default swaps—could further test the model’s generalizability.
-Additionally, exploring different time intervals, such as hourly, minute, or even second-level data, could provide deeper insights, as this study exclusively utilized daily intervals.
-Testing the model on illiquid assets could further validate its robustness in different market conditions.
-In addition, incorporating alternative data sources, such as social media sentiment, news feeds, or satellite imagery, might further enhance model accuracy.
-Finally, future research could focus on optimizing the LSTM-based model for real-world applications, such as integrating it into real-time trading systems or risk management frameworks.
-Statements and Declarations
-The authors declare that there are no competing interests related to this work.
-Appendix A Financial Assets
+This paper introduces a novel approach to estimating the distribution of log returns by combining quantile regression with smoothed density estimation, providing a robust framework for capturing tail risks and distributional characteristics across diverse asset classes.
+By leveraging asset-neutral features and general market variables, our models are applicable to a wide range of financial datasets, making them versatile tools for risk management, portfolio optimization, and derivative pricing.
+The proposed qLSTM model consistently demonstrated superior performance on real-world data, achieving the lowest quantile loss across most asset classes, while the qHybrid model excelled in synthetic datasets, particularly for skewed or heavy-tailed distributions.
+These results underscore the strengths of both models in addressing the limitations of traditional distributional assumptions, such as normality, in financial returns.
+The evaluation metrics, including Wasserstein distance, CRPS, and VaR, highlight the complementary nature of the qLSTM and qHybrid models.
+The qLSTM model excelled in extreme tail risk estimation, capturing critical distributional behaviors for volatile asset classes like cryptocurrencies and equity indices.
+Meanwhile, the qHybrid model provided improved calibration for broader confidence levels and distributions, demonstrating its adaptability to various market conditions.
+However, some variability in performance across asset classes suggests that no single model is universally optimal, emphasizing the need for tailored approaches based on specific data characteristics.
+To enhance the literature, the paper allows for reproducibility by publicly hosting the code used for this study on GitHub.
+Furthermore, all data is sourced via an open API and the script to download the required data can be found in the GitHub repository.
+This makes comparisons and iterations on the works proposed, more accessible thereby fostering research robustness.
+Future research can explore other methods to create better Hybrid models, particularly by leveraging certain quantiles with dynamic weighting schemes to optimize the combination of Gaussian and model-predicted quantiles.
+This could provide a more adaptable and context-specific qHybrid approach to estimating distributions.
+Moreover, extending the model to handle ultra-high-frequency data or incorporating more sophisticated market features, such as macroeconomic indicators or sentiment analysis, could further enhance its applicability.
+These avenues represent exciting opportunities to refine and expand the utility of distributional modeling in finance.
+Appendix A Quantiles
+The quantiles used in this study are:
+[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03,[ 0.00005 , 0.00025 , 0.00075 , 0.00125 , 0.00175 , 0.0025 , 0.005 , 0.01 , 0.015 , 0.02 , 0.03 ,
+$0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7,$
+0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995 0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995]0.75 , 0.8 , 0.85 , 0.9 , 0.95 , 0.98 , 0.99 , 0.995 , 0.9975 , 0.99925 , 0.99975 , 0.99995 ]
+Appendix B Quantile-Smooth Density Estimation
+This method generates a smooth probability density function (PDF) from a set of quantiles and their associated probabilities, ensuring that the resulting representation adheres to the properties of a valid probability distribution.
+Starting with the quantiles, it removes duplicate or nearly identical points to maintain numerical stability, then creates a denser grid for better resolution.
+Using a monotonic cubic spline or fallback interpolation, it estimates the CDF, which is clamped to the range [0, 1] and adjusted to be strictly non-decreasing.
+The method rescales the CDF to start at 0 and end at 1, guaranteeing its validity as a cumulative function.
+To derive the PDF, the method computes the CDF gradient, applying smoothing techniques such as convolution to remove spikes and ensure a well-behaved density curve.
+Non-negativity is enforced, and the PDF is normalized to integrate to one.
+The final output includes the smoothed PDF, the regenerated CDF, and the grid over which they are defined, offering a high-quality approximation of the distribution.
+Optional visualization helps validate the results, showing both the smoothed PDF and the corresponding CDF with respect to the input quantiles.
+This makes it a robust method for converting quantiles into a smooth and interpretable density function suitable for further analysis or modeling.
+The function is based on three priors, $\epsilon$ which controls how close different quantiles are allowed to each other, min density which controls the minimum amount of density allowed in the PDF, and window which controls how many extreme the convolutions are for smoothing.
+The optimal priors are found via TPE on the validation data, similar to how the hyperparameters are found for the models themselves.
+Table 10 shows the best values found during optimization.
+Model $\epsilon$ min density window
+qLSTM 0.000253 0.027341 113
+hybrid 0.000899 0.033704 43
+Table 10: Optimal priors for the smooth PDF approximation
+Below is the Python code used for the Quantile-Smooth Density Estimation conversion in this work.
+Listing 1: Quantiles to PDF using Quantile-Smooth Density Estimation
+def generate_smooth_pdf(quantiles, taus, min_density=1e-3, eps=1e-6, window=61):
+␣␣␣␣GenerateasmoothedPDFfromquantileswithadditionalcontrolstopreventspikes
+␣␣␣␣andensuretheCDFisbetween[0,1].
+# Constants
+GRID_POINTS = 1000
+og_quants = quantiles.copy()
+og_taus = taus.copy()
+unique_mask = np.concatenate(([True], np.diff(quantiles) > eps))
+quantiles = quantiles[unique_mask]
+taus = taus[unique_mask]
+# Create denser grid
+grid_x = np.linspace(quantiles[0], quantiles[-1], GRID_POINTS)
+# Monotonic spline for the CDF
+try:
+cdf_monotonic = PchipInterpolator(quantiles, taus, extrapolate=False)
+cdf = cdf_monotonic(grid_x)
+except Exception as e:
+cdf = np.interp(grid_x, quantiles, taus)
+# Clamp CDF to [0,1], then ensure it’s monotonically non-decreasing
+cdf = np.clip(cdf, 0, 1)
+cdf = np.maximum.accumulate(cdf)
+# Rescale so that it starts exactly at 0 and ends exactly at 1
+cdf -= cdf[0]
+if cdf[-1] > 0:
+cdf /= cdf[-1]
+# Approximate PDF from finite differences (or use derivative if PCHIP)
+density = np.gradient(cdf, grid_x)
+smoothed_density = np.convolve(density, np.ones(window)/window, mode=’same’)
+# Ensure non-negative and non-zero density
+smoothed_density = np.maximum(smoothed_density, min_density)
+# Normalize PDF to integrate to 1
+area = np.trapz(smoothed_density, grid_x)
+smoothed_density = smoothed_density / area
+# regenerate CDF
+cdf = np.cumsum(smoothed_density) * (grid_x[1] - grid_x[0])
+return grid_x, smoothed_density, cdf
+Appendix C Data
+The following sections detail the data used for this study.
+C.1 Financial Assets
 The table below provides a comprehensive list of financial assets, including their ticker symbols, associated indexes, start dates, annualized mean returns, and standard deviations of returns, rounded to two decimal places where applicable.
@@
 Orange Juice Futures JO=F Commodities 2001-10-17 0.37 33.20
-Bitcoin BTC-USD Cryptocurrencies 2014-10-09 2.25 59.09
-Ethereum ETH-USD Cryptocurrencies 2017-12-01 1.16 76.15
-Ripple XRP-USD Cryptocurrencies 2017-12-01 0.64 96.11
-Cardano ADA-USD Cryptocurrencies 2017-12-01 1.14 94.15
-Dogecoin DOGE-USD Cryptocurrencies 2017-12-01 2.69 115.10
-Binance Coin BNB-USD Cryptocurrencies 2017-12-01 3.61 86.33
-NEAR Protocol NEAR-USD Cryptocurrencies 2020-11-05 2.65 109.42
-Bitcoin Cash BCH-USD Cryptocurrencies 2017-12-01 -1.20 93.81
-Litecoin LTC-USD Cryptocurrencies 2014-10-09 1.38 85.60
-Solana SOL-USD Cryptocurrencies 2020-05-02 5.85 114.08
-Appendix B Market Data
+Bitcoin BTC-USD Crypto 2014-10-09 2.25 59.09
+Ethereum ETH-USD Crypto 2017-12-01 1.16 76.15
+Ripple XRP-USD Crypto 2017-12-01 0.64 96.11
+Cardano ADA-USD Crypto 2017-12-01 1.14 94.15
+Dogecoin DOGE-USD Crypto-currencies 2017-12-01 2.69 115.10
+Binance Coin BNB-USD Crypto-currencies 2017-12-01 3.61 86.33
+NEAR Protocol NEAR-USD Crypto-currencies 2020-11-05 2.65 109.42
+Bitcoin Cash BCH-USD Crypto-currencies 2017-12-01 -1.20 93.81
+Litecoin LTC-USD Crypto-currencies 2014-10-09 1.38 85.60
+Solana SOL-USD Crypto-currencies 2020-05-02 5.85 114.08
+C.2 Market Data
 The table below lists the market data used in this study, including tickers, a brief representation, start dates, annualized mean returns, and standard deviations.
@@
 Corporate Bond Index AGG U.S. investment-grade bond market performance 2000-02-03 0.20 5.75
-Appendix C qDense Training Details & Results
-The following tables summarize the hyperparameters used during the training of the qDense model.
-The first table presents the ranges explored during the hyperparameter tuning process, and the second table lists the final hyperparameters used in the model, rounded down to the fourth decimal place.
-Table 9: Hyperparameters explored for qDense Asset-Specific and Market Data Models
-Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
-Batch size $b\in\{32,64,128,256,512,1024,2048\}$
-Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
-Normalization window $w\in[5,250]$
-Dropout rate $D\in[0.0,0.9]$
-L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
-L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
-Use layer normalization Boolean: $\{1,0\}$
-Raw hidden layers $l_{\text{raw, asset}}\in[1,5]$ $l_{\text{raw, market}}\in[1,5]$
-Hidden layer units $u_{\text{raw, asset}}\in\{16,\cdots,512\}$ $u_{\text{raw, market}}\in\{16,\cdots,512\}$
-[…]
-Table 10: Final Hyperparameters for the qDense model, rounded down to the 4th decimal place
-Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
-Batch size 2048
-Learning rate 0.0007
-Normalization window 123
-Raw hidden layers 2 4
-Hidden layer units 64, 512 64, 512, 16, 32
-Dropout rate 0.1780
-Market activation Sigmoid
-Hidden activation ELU
-Use layer normalization Yes (1)
-L1 regularization 0.0007
-L2 regularization 0.0004
-Appendix D Kernel Density Estimation
-Kernel Density Estimation (KDE) is a non-parametric method used to estimate the probability density function (PDF) of a random variable.
-In this study, KDE was used to convert predicted quantiles into a smooth probability density function.
-The advantage of KDE is that it allows for flexible modeling of the underlying data distribution without assuming any particular parametric form, which is useful when working with financial return distributions that may not follow standard distributions like normal or log-normal.
-The quantiles predicted by the model are transformed into a smooth PDF using KDE, ensuring that the final distribution is continuous and appropriately represents the data.
-This process helps provide more accurate density estimates, which are crucial for downstream tasks such as risk estimation and derivative pricing.
-Below is the Python code used for the KDE conversion in this work.
-1import numpy as np
-2import matplotlib.pyplot as plt
-3from sklearn.neighbors import KernelDensity
-4
-5def quantiles_to_pdf_kde(quantiles, grid_points=200, bandwidth=0.01):
-6 """
-7 Convert quantiles to a probability density function (PDF) using Kernel Density Estimation (KDE).
-8
-9 Parameters:
-10 - quantiles: Array of quantile values.
-11 - grid_points: Number of grid points for a denser representation.
-12 - bandwidth: Bandwidth for the KDE (controls the smoothness of the density).
-13
-14 Returns:
-15 - pdf_grid: Probability density function values on the dense grid.
-16 - cdf_grid: Cumulative distribution function values on the dense grid (obtained by integrating the PDF).
-17 - dense_grid: The dense grid of quantiles corresponding to the PDF and CDF values.
-18 """
-19
-20 # Create a denser grid between the quantiles
-21 dense_grid = np.linspace(min(quantiles), max(quantiles), grid_points).reshape(-1, 1)
-22
-23 # Fit KDE to the quantile values
-24 kde = KernelDensity(kernel=’gaussian’, bandwidth=bandwidth).fit(quantiles.reshape(-1, 1))
-25
-26 # Compute log density and convert to regular density
-27 log_dens = kde.score_samples(dense_grid)
-28 pdf_grid = np.exp(log_dens)
-29
-30 # Approximate the CDF by cumulatively summing the PDF
-31 cdf_grid = np.cumsum(pdf_grid)
-32 cdf_grid /= cdf_grid[-1] # Normalize to [0, 1]
-33
-34 return pdf_grid, cdf_grid, dense_grid
-Listing 1: Quantiles to PDF using KDE
-Appendix E Synthetic Data
+C.3 Synthetic Data
 In this study, synthetic data was generated across four different distributions: Normal, Log-Normal, Gamma, and Uniform.
@@
 Table 13: Generated $\mu$ and $\sigma$ values and correlation with market data across different test datasets.
-Distribution Type Asset Generated $\mu$, $\sigma$ Correlation with Market Data
+Dist Type Asset Generated $\mu$, $\sigma$ Correl -ation with Market Data
 Normal Asset 1 $\mu=-0.000663$, $\sigma=0.022852$ 0.7047
@@
 Asset 10 $\mu=-0.000876$, $\sigma=0.012219$ 0.8545
+[h]
 In addition to the statistical properties shown in the table, the synthetic data was generated using the following Python function.
 The function allows for the generation of log returns that are correlated with real market data by controlling the mean ($\mu$) and standard deviation ($\sigma$) of the generated data.
-1import numpy as np
-2import pandas as pd
-3from scipy.stats import gamma, lognorm, uniform
-4
-5def generate_synthetic_data_with_target_correlation(market_data, target_correlation, mu, sigma, distribution=’normal’):
-6 """
-7 Generate a synthetic dataset with returns that have a target correlation to the market_data.
-8
-9 Parameters:
-10 - market_data: numpy array of log returns from the market data
-11 - target_correlation: the target correlation for synthetic data’s log returns to the market data
-12 - mu: mean of the synthetic dataset
-13 - sigma: standard deviation of the synthetic dataset
-14 - distribution: the distribution to sample synthetic data from (’normal’, ’gamma’, ’lognormal’, ’uniform’)
-15
-16 Returns:
-17 - df: DataFrame with synthetic prices
-18 """
-19
-20 num_periods = len(market_data)
-21
-22 # Step 1: Generate independent noise based on the specified distribution
-23 if distribution == ’normal’:
-24 noise = np.random.normal(loc=0, scale=1, size=num_periods) # Standard normal noise
-25 elif distribution == ’gamma’:
-26 shape = 2 # Shape parameter for the gamma distribution
-27 noise = gamma.rvs(a=shape, scale=1, size=num_periods)
-28 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
-29 elif distribution == ’lognormal’:
-30 noise = lognorm.rvs(s=1, scale=np.exp(0), size=num_periods)
-31 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
-32 elif distribution == ’uniform’:
-33 noise = uniform.rvs(loc=-1, scale=2, size=num_periods) # Uniform distribution [-1, 1]
-34 else:
-35 raise ValueError("Unsupported distribution. Choose from ’normal’, ’gamma’, ’lognormal’, or ’uniform’.")
-36
-37 # Step 2: Standardize market data
-38 market_std = (market_data - np.mean(market_data)) / np.std(market_data)
-39
-40 # Step 3: Create a linear combination of market data and noise for correlated returns
-41 synthetic_log_returns = target_correlation * market_std + np.sqrt(1 - target_correlation**2) * noise
-42
-43 # Step 4: Adjust the synthetic returns to have the desired mean and standard deviation
-44 synthetic_log_returns = synthetic_log_returns * sigma + mu
-45
-46 # Step 5: Convert synthetic log returns to price data
-47 initial_price = np.random.uniform(5, 1000) # Random initial price
-48 synthetic_prices = initial_price * np.exp(np.cumsum(synthetic_log_returns))
-49
-50 # Step 6: Create a DataFrame with synthetic price data
-51 df = pd.DataFrame(synthetic_prices, columns=[’Adj Close’])
-52
-53 # Add High, Low, Volume columns to allow all functions to work
-54 df[’High’] = df[’Adj Close’] * (1 + np.random.normal(0, 0.01, size=len(df)))
-55 df[’Low’] = df[’Adj Close’] * (1 - np.random.normal(0, 0.01, size=len(df)))
-56 df[’Volume’] = 0
-57
-58 df = generate_technical_features(df)
-59
-60 return df
 Listing 2: Synthetic Data Generation with Target Correlation
+import numpy as np
+import pandas as pd
+from scipy.stats
+import gamma, lognorm, uniform
+def generate_synthetic_data_with_target_correlation(market_data, target_correlation, mu, sigma, distribution=’normal’):
+␣␣␣␣Generateasyntheticdatasetwithreturnsthathaveatargetcorrelationtothemarket_data.
+␣␣␣␣Parameters:
+␣␣␣␣-market_data:numpyarrayoflogreturnsfromthemarketdata
+␣␣␣␣-target_correlation:thetargetcorrelationforsyntheticdata’slogreturnstothemarketdata
+␣␣␣␣-mu:meanofthesyntheticdataset
+␣␣␣␣-sigma:standarddeviationofthesyntheticdataset
+␣␣␣␣-distribution:thedistributiontosamplesyntheticdatafrom(’normal’,’gamma’,’lognormal’,’uniform’)
+␣␣␣␣Returns:
+␣␣␣␣-df:DataFramewithsyntheticprices
+num_periods = len(market_data)
+# Step 1: Generate independent noise based on the specified distribution
+if distribution == ’normal’:
+noise = np.random.normal(loc=0, scale=1, size=num_periods) # Standard normal noise
+elif distribution == ’gamma’:
+shape = 2 # Shape parameter for the gamma distribution
+noise = gamma.rvs(a=shape, scale=1, size=num_periods)
+noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
+elif distribution == ’lognormal’:
+noise = lognorm.rvs(s=1, scale=np.exp(0), size=num_periods)
+noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
+elif distribution == ’uniform’:
+noise = uniform.rvs(loc=-1, scale=2, size=num_periods) # Uniform distribution [-1, 1]
+else:
+raise ValueError(”Unsupporteddistribution.Choosefrom’normal’,’gamma’,’lognormal’,or’uniform’.”)
+# Step 2: Standardize market data
+market_std = (market_data - np.mean(market_data)) / np.std(market_data)
+# Step 3: Create a linear combination of market data and noise for correlated returns
+synthetic_log_returns = target_correlation * market_std + np.sqrt(1 - target_correlation**2) * noise
+# Step 4: Adjust the synthetic returns to have the desired mean and standard deviation
+synthetic_log_returns = synthetic_log_returns * sigma + mu
+# Step 5: Convert synthetic log returns to price data
+initial_price = np.random.uniform(5, 1000) # Random initial price
+synthetic_prices = initial_price * np.exp(np.cumsum(synthetic_log_returns))
+# Step 6: Create a DataFrame with synthetic price data
+df = pd.DataFrame(synthetic_prices, columns=[’AdjClose’])
+# Add High, Low, Volume columns to allow all functions to work
+df[’High’] = df[’AdjClose’] * (1 + np.random.normal(0, 0.01, size=len(df)))
+df[’Low’] = df[’AdjClose’] * (1 - np.random.normal(0, 0.01, size=len(df)))
+df[’Volume’] = 0
+df = generate_technical_features(df)
+return df
 The function generates synthetic log returns that are then converted to price data, while ensuring a target correlation with the real market data.
 This allows for a more controlled and realistic evaluation of the model’s performance on different types of financial data.
-Appendix F Quantiles
-The quantiles used in this study are:
-[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03[0.00005,0.00025,0.00075,0.00125,0.00175,0.0025,0.005,0.01,0.015,0.02,0.03,[ 0.00005 , 0.00025 , 0.00075 , 0.00125 , 0.00175 , 0.0025 , 0.005 , 0.01 , 0.015 , 0.02 , 0.03 ,
-$0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7,$
-0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995 0.75,0.8,0.85,0.9,0.95,0.98,0.99,0.995,0.9975,0.99925,0.99975,0.99995]0.75 , 0.8 , 0.85 , 0.9 , 0.95 , 0.98 , 0.99 , 0.995 , 0.9975 , 0.99925 , 0.99975 , 0.99995 ]
+Appendix D Training Details
+The following sections give details on training procedures including both training of the models as well as optimal hyper parameters.
+D.1 qLSTM Hyperparamters
+The following table shows the optimal hyperparameters for the qLSTM model.
+Table 14: Final Hyperparameters for the qLSTM model, rounded down to the 4th decimal place
+Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
+Batch size 64
+Learning rate 0.0006
+Normalization window 219
+LSTM layers 1 1
+LSTM units 16 16
+Dense layers 5 3
+Dense units per layer 128, 64, 64, 32, 32 16, 16, 32
+Dropout rate 0.1778
+Market activation Tanh
+Hidden activation ELU
+Use layer normalization no (0)
+L1 regularization 0.0006
+L2 regularization 0.0009
+D.2 qDense Training Details & Results
+The following tables summarize the hyperparameters used during the training of the qDense model.
+The first table presents the ranges explored during the hyperparameter tuning process, and the second table lists the final hyperparameters used in the model, rounded down to the fourth decimal place.
+Table 15: Hyperparameters explored for qDense Asset-Specific and Market Data Models
+Hyper-parameter Asset Model (Stage 1) Market Model (Stage 2)
+Batch size $b\in\{32,64,128,256,512,1024,2048\}$
+Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
+Normalization window $w\in[5,250]$
+Dropout rate $D\in[0.0,0.9]$
+L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
+L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
+Use layer normalization Boolean: $\{1,0\}$
+Raw hidden layers $l_{\text{raw, asset}}\in[1,5]$ $l_{\text{raw, market}}\in[1,5]$
+Hidden layer units $u_{\text{raw, asset}}\in\{16,\cdots,512\}$ $u_{\text{raw, market}}\in\{16,\cdots,512\}$
+[…]
+Table 16: Final Hyperparameters for the qDense model, rounded down to the 4th decimal place
+Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
+Batch size 32
+Learning rate 0.0003
+Normalization window 207
+Raw hidden layers 3 1
+Hidden layer units 32, 32, 64 64
+Dropout rate 0.4623
+Market activation ELU
+Hidden activation Leaky RELU
+Use layer normalization No (0)
+L1 regularization 0.0005
+L2 regularization 0.0001